### v1

# Kuznecov formulae for fractal measures

Abstract.
Let $(M,g)$ be a compact, connected Riemannian manifold of dimension $n\geq 2$, and let $\{e_{j}\}_{j=0}^{\infty}$ be an orthonormal basis of Laplace eigenfunctions $\Delta_{g}e_{j}=\lambda_{j}^{2}e_{j}$. Given a finite Borel measure $\mu$ on $M$, consider the Kuznecov sum
$N_{\mu}(\lambda):=\sum_{\lambda_{j}\leq\lambda}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}.$
Assume that $\mu$ is $s$-Ahlfors regular for some $s\in(0,n)$ and admits an averaged $s$-density constant $A_{\mu}$. We prove that
$N_{\mu}(\lambda)=(2\pi)^{-(n-s)}\,\operatorname{vol}(B^{\,n-s})\,A_{\mu}\,\lambda^{n-s}+o(\lambda^{n-s})\qquad(\lambda\to\infty).$
The hypotheses of $s$-Ahlfors regularity and the averaged $s$-density condition are essentially optimal for such a one-term asymptotic, and in general the remainder $o(\lambda^{n-s})$ cannot be improved uniformly to a power-saving error term. This extends the classical Kuznecov formulae of Zelditch [Zel92] for smooth submanifold measures to a broad class of singular and fractal measures.
1. Introduction
Let $(M,g)$ be a compact, connected, smooth Riemannian manifold of dimension $n\geq 2$, and let
$\Delta_{g}e_{j}=\lambda_{j}^{2}e_{j},\qquad 0=\lambda_{0}<\lambda_{1}\leq\lambda_{2}\leq\cdots,$
denote a complete orthonormal basis of real Laplace eigenfunctions. Given a finite Borel measure $\mu$ on $M$, we consider the associated Kuznecov sum
(1.1) $N_{\mu}(\lambda):=\sum_{\lambda_{j}\leq\lambda}\Big|\int_{M}e_{j}\,d\mu\Big|^{2},\qquad\lambda\geq 1.$
A basic and well-studied case is when $\mu$ is the Riemannian volume measure on an embedded submanifold. Let $H\subset M$ be an embedded submanifold of dimension $k$ with induced Riemannian metric and volume measure $dV_{H}$. In his seminal work [Zel92], Zelditch proved a Kuznecov sum formula for these coefficients, showing that
(1.2) $\sum_{\lambda_{j}\leq\lambda}\biggl|\int_{H}e_{j}\,dV_{H}\biggr|^{2}=(2\pi)^{-(n-k)}\,\operatorname{vol}(B^{n-k})\,\operatorname{vol}(H)\,\lambda^{\,n-k}+O(\lambda^{n-k-1}),$
where $B^{n-k}\subset\mathbb{R}^{n-k}$ is the Euclidean unit ball and $\operatorname{vol}(H)$ denotes the kk–dimensional Riemannian volume of $H$. Thus the leading growth of the Kuznecov sum is governed by the ambient dimension $n$, the submanifold dimension $k$, and an explicit geometric factor.
More recently, Wyman and the author [WX23b] refined Zelditch’s formula by obtaining a two-term asymptotic for (1.2), identifying the next-order oscillatory term. In these works the measure is always smooth and supported on a submanifold. On the other hand, there has been growing recent interest in the behavior of eigenfunctions restricted to general Borel sets. For example, Eswarathasan and Pramanik [EP22] initiated a program to study the $L^{p}$ norms of eigenfunctions with respect to a fractal measure $\mu$, connecting classical results of Sogge [Sog88] and Burq–Gérard–Tzvetkov [BGT07]. An essentially sharp version of such bounds in two dimensions was recently obtained by Gao, Miao, and the author [GMX24]. The goal of this paper is to extend Zelditch’s Kuznecov formulae to a much broader class of measures, including singular and fractal measures, while retaining the same leading term.
Throughout the paper we use the following convention for Euclidean ball volumes. For $m>0$ (not necessarily an integer), we write $\operatorname{vol}(B^{m})$ for the “mm–dimensional volume” of the Euclidean unit ball in formal dimension $m$, defined by
$\operatorname{vol}(B^{m}):=\frac{\pi^{m/2}}{\Gamma\bigl(\frac{m}{2}+1\bigr)}.$
When $m\in\mathbb{N}$, this coincides with the usual Lebesgue mm–volume of the unit ball in $\mathbb{R}^{m}$. For an embedded kk–dimensional submanifold $H\subset M$ we write $\operatorname{vol}(H)$ for its kk–dimensional Riemannian volume.
We now describe the small-scale geometric hypotheses on $\mu$ that will play the role of “dimension” and “density” in our setting. Throughout, $B(x,r)$ denotes the geodesic ball of radius $r$ centered at $x\in M$ with respect to the Riemannian distance $d_{g}$.
Definition 1.1 ($s$-Ahlfors regular measure).
Let $s\in(0,n)$. A finite Borel measure $\mu$ on $M$ is called $s$-Ahlfors regular if there exist constants $c,C>0$ and $r_{0}>0$ such that
(1.3) $c\,r^{s}\leq\mu(B(x,r))\leq C\,r^{s},\qquad x\in\operatorname{supp}\mu,\ 0<r\leq r_{0}.$
This condition is standard in geometric measure theory and implies that the Hausdorff dimension of $\mu$ is $s$. See, for example, [Mat15]. It is satisfied by Riemannian volume measures on $s$-dimensional embedded submanifolds, as well as by many self-similar and self-conformal measures on fractals.
To obtain a one-term asymptotic for (1.1), one needs a stabilization of the averaged small-scale behavior of $\mu$. For $r\geq 0$ define the distance distribution
(1.4) $F(r):=(\mu\times\mu)\bigl(\{(x,y)\in M\times M:\ d_{g}(x,y)\leq r\}\bigr)=\int_{M}\mu\bigl(B(x,r)\bigr)\,d\mu(x).$
Definition 1.2 (Averaged $s$-density).
Let $s\in(0,n)$. We say that $\mu$ admits an averaged $s$-density if there exists a constant $A_{\mu}\in(0,\infty)$ (called the averaged $s$-density constant of $\mu$) such that
(1.5) $F(r)=\operatorname{vol}(B^{s})\,A_{\mu}\,r^{s}+o(r^{s})\qquad(r\to 0^{+}).$
Equivalently,
(1.6) $A_{\mu}=\lim_{r\to 0^{+}}\frac{F(r)}{\operatorname{vol}(B^{s})\,r^{s}}=\lim_{r\to 0^{+}}\frac{1}{\operatorname{vol}(B^{s})\,r^{s}}\int_{M}\mu(B(x,r))\,d\mu(x).$
We can now state our main result. It shows that under the Ahlfors regularity assumption and the averaged $s$-density condition (1.5), the Kuznecov sum (1.1) has a one-term asymptotic with exponent nsn-s and an explicit leading constant depending only on $n$ and $s$.
Theorem 1.3.
Let $(M,g)$ be a compact, connected Riemannian manifold of dimension $n\geq 2$. Fix $s\in(0,n)$ and let $\mu$ be a finite Borel measure on $M$ which is $s$-Ahlfors regular in the sense of (1.3) and admits an averaged $s$-density constant $A_{\mu}$ in the sense of (1.5). Then the Kuznecov sum (1.1) satisfies
(1.7) $N_{\mu}(\lambda)=C_{n,s}\,A_{\mu}\,\lambda^{\,n-s}+o(\lambda^{\,n-s})\qquad(\lambda\to\infty),$
where
$C_{n,s}=(2\pi)^{-(n-s)}\,\operatorname{vol}(B^{\,n-s}).$
The proof combines a heat-kernel regularization of the Kuznecov sum with a Tauberian argument, with the leading term determined by the small-scale distance distribution of $\mu$.
Remark 1.4.
The averaged $s$-density condition (1.5) is automatically satisfied under the stronger pointwise density hypothesis. Namely, suppose that $\mu$ is $s$-Ahlfors regular and admits a pointwise $s$-density $\theta$ in the sense that
$\theta(x)=\lim_{r\to 0^{+}}\frac{\mu(B(x,r))}{\operatorname{vol}(B^{s})\,r^{s}}$
exists for $\mu$-a.e. $x$. Then by the Ahlfors upper bound $\mu(B(x,r))/(\operatorname{vol}(B^{s})r^{s})\lesssim 1$ and dominated convergence,
$\frac{F(r)}{\operatorname{vol}(B^{s})\,r^{s}}=\int_{M}\frac{\mu(B(x,r))}{\operatorname{vol}(B^{s})\,r^{s}}\,d\mu(x)\longrightarrow\int_{M}\theta(x)\,d\mu(x),\qquad(r\to 0^{+}).$
Hence (1.5) holds with
$A_{\mu}=\int_{M}\theta\,d\mu.$
In particular, (1.5) is strictly weaker than the existence of a pointwise $s$-density.
The leading exponent in (1.7) is nsn-s, determined by the ambient dimension $n$ and the Hausdorff dimension $s$ of $\mu$, and the leading coefficient factors as a universal constant times the averaged $s$-density constant $A_{\mu}$. In the special case where $\mu=dV_{H}$ is the Riemannian volume measure on an embedded submanifold $H$ of dimension $s$, $\mu$ admits a pointwise $s$-density $\theta$ and hence satisfies the averaged $s$-density condition (1.5) by Remark 1.4. In fact $\theta\equiv 1$ and $A_{\mu}=\operatorname{vol}(H)$, so (1.7) reduces to
$\sum_{\lambda_{j}\leq\lambda}\biggl|\int_{H}e_{j}\,dV_{H}\biggr|^{2}=(2\pi)^{-(n-s)}\,\operatorname{vol}(B^{\,n-s})\,\operatorname{vol}(H)\,\lambda^{\,n-s}+o(\lambda^{\,n-s}),$
whose leading term agrees with Zelditch’s Kuznecov formulae (1.2) with $k=s$. Thus Theorem 1.3 extends the classical Kuznecov formulae to a large class of singular and fractal measures. To the best of our knowledge, Theorem 1.3 is the first Kuznecov-type asymptotic for fractal measures.
Moreover, as discussed in [WX23a], the left-hand side of Zelditch’s Kuznecov formula (1.2) can be viewed as “audible” spectral information associated with the submanifold $H$, while the right-hand side encodes geometric information about $H$, namely its dimension $k$ and its $k$-dimensional Riemannian volume $\operatorname{vol}(H)$. Theorem 1.3 extends this audible–geometric correspondence from smooth submanifold measures to $s$-Ahlfors regular measures admitting an averaged $s$-density constant, so that the Hausdorff dimension $s$ and the effective “volume” $A_{\mu}$ of the measure are also audible through the leading asymptotics of $N_{\mu}(\lambda)$.
For $0<u<n$ let
(1.8) $I_{u}(\mu):=\iint d_{g}(x,y)^{-u}\,d\mu(x)\,d\mu(y)$
denote the Riesz energy of $\mu$ with respect to the Riemannian distance $d_{g}(x,y)$ on $M$. There is an earlier work of Hare and Roginskaya [HR03] which is closely related in spirit to our results. They established energy formulae on compact Riemannian manifolds which relate $I_{u}(\mu)$ to weighted sums of the projections of $\mu$ onto Laplace eigenspaces. In Section 4 we show that their identities imply the following general upper bound without assuming the existence of an averaged $s$-density constant.
Theorem 1.5.
Let $(M,g)$ be a compact Riemannian manifold of dimension $n\geq 2$, and let $\mu$ be a finite Borel measure on $M$. Fix $u\in(0,n)$ and suppose that the Riesz energy $I_{u}(\mu)$ defined in (1.8) is finite. Then there exists a constant $C_{1}=C_{1}(M,g,u,\mu)$ such that
$\sum_{\lambda_{j}\leq\lambda}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}\leq C_{1}\,\lambda^{\,n-u}\qquad(\lambda\geq 1).$
In particular, if $\mu$ is $s$-Ahlfors regular for some $s\in(0,n)$, then $I_{u}(\mu)<\infty$ for every $u<s$, and hence for each $\varepsilon>0$ there exists a constant $C_{2}=C_{2}(M,g,s,\varepsilon,\mu)$ such that
$\sum_{\lambda_{j}\leq\lambda}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}\leq C_{2}\,\lambda^{\,n-s+\varepsilon}\qquad(\lambda\geq 1).$
Thus, for measures with finite Riesz energy, one also has nontrivial polynomial upper bounds on the Kuznecov sum. However, in the formulae of Hare–Roginskaya the exponent $u$ is fixed, and the constants appearing in Theorem 1.5 are not controlled uniformly as $u\to s^{-}$. Since our proof of Theorem 1.3 requires a precise description of the singular behavior of $I_{u}(\mu)$ at $u=s$ in order to recover the leading constant in (1.7), one cannot simply invoke their identities and pass to the limit $u\to s$.
Sharpness
We point out that the assumptions in Theorem 1.3 and Theorem 1.5 are essentially optimal. In Section 5 we give examples showing, on the one hand, that one cannot dispense with the $s$-Ahlfors regularity assumption if one wishes to control the exponent purely in terms of the Hausdorff dimension of $\mu$, and on the other hand, that some stabilization of the small-scale distance distribution—such as the averaged $s$-density condition (1.5)—is needed to obtain a genuine one-term asymptotic. Under $s$-Ahlfors regularity alone one can at best expect upper and lower bounds of order $\lambda^{n-s}$, but not the existence of a limit for $N_{\mu}(\lambda)/\lambda^{n-s}$.
In addition to these structural restrictions, we show in Section 5 that the remainder term in Theorem 1.3 cannot, in general, be improved to a uniform power saving over the class of measures under consideration. More precisely, we construct explicit examples showing that there is no exponent $\delta>0$ such that for every $s$-Ahlfors regular measure with an averaged $s$-density constant $A_{\mu}$ one has
$N_{\mu}(\lambda)=C_{n,s}A_{\mu}\,\lambda^{n-s}+O\bigl(\lambda^{n-s-\delta}\bigr),$
with an implicit constant depending only on $(M,g,n,s,\mu)$. The $o(\lambda^{n-s})$ remainder in (1.7) is therefore essentially sharp at this level of generality.
Organization
In Section 2 we introduce a heat-kernel regularization of the Kuznecov sum, and establish its small-time asymptotics in terms of the averaged $s$-density constant $A_{\mu}$. Section 3 interprets this smoothed Kuznecov sum as a Laplace–Stieltjes transform and applies a Tauberian theorem to prove Theorem 1.3 and identify the constant $C_{n,s}$. In Section 4 we revisit the energy identities of Hare and Roginskaya to derive general polynomial upper bounds for $N_{\mu}(\lambda)$, proving Theorem 1.5. Finally, Section 5 is devoted to sharpness: we give examples showing that the Ahlfors regularity and averaged $s$-density hypotheses in Theorem 1.3 are essentially optimal, and that the $o(\lambda^{n-s})$ remainder cannot, in general, be improved to a uniform power-saving error term.
Acknowledgement
This project is supported by the National Key R&D Program of China under Grant No. 2022YFA1007200, the Natural Science Foundation of China under Grant No. 12571107, and the Zhejiang Provincial Natural Science Foundation of China under Grant No. LR25A010001. The author thanks Xucheng Hu and Diankun Liu for carefully reading an earlier draft of this paper.
Notation
We write $A\lesssim B$ to mean that there exists an implicit constant $C>0$ such that
$A\leq C\,B.$
We write $A\gtrsim B$ if $B\lesssim A$, and $A\approx B$ if both $A\lesssim B$ and $B\lesssim A$.
For nonnegative functions $f,g$ and $x_{0}\in\mathbb{R}\cup\{\infty\}$ we write
$f(x)\sim g(x)\qquad(x\to x_{0})$
to mean
$\lim_{x\to x_{0}}\frac{f(x)}{g(x)}=1.$
2. Heat kernel and a smoothed Kuznecov sum
In this section, we introduce a heat-kernel-regularized Kuznecov sum and identify its small-time behavior in terms of the averaged $s$-density constant $A_{\mu}$.
2.1. Heat kernel asymptotics near the diagonal
Throughout we write $p_{t}(x,y)$ for the heat kernel on $(M,g)$:
$(\partial_{t}-\Delta_{g})p_{t}(x,y)=0,\qquad\lim_{t\to 0^{+}}p_{t}(\,\cdot\,,y)=\delta_{y},$
so that for every $f\in C^{\infty}(M)$,
$e^{t\Delta_{g}}f(x)=\int_{M}p_{t}(x,y)f(y)\,dV(y)=\sum_{j=0}^{\infty}e^{-t\lambda_{j}^{2}}\langle f,e_{j}\rangle e_{j}(x).$
The small-time behavior of $p_{t}$ near the diagonal is classical. See, for example, [BGV92, Ch. 2] or [Gri09, Ch. 8].
Lemma 2.1.
There exist $r_{0}>0$, $t_{0}>0$, and smooth functions $u_{k}(x,y)$ on $\{(x,y)\in M\times M:d_{g}(x,y)<2r_{0}\}$, $k=0,1,2,\dots$, with $u_{0}(x,x)\equiv 1$, such that for every $N\in\mathbb{N}$ we have
(2.1) $p_{t}(x,y)=(4\pi t)^{-n/2}e^{-d_{g}(x,y)^{2}/(4t)}\biggl(\sum_{k=0}^{N-1}t^{k}u_{k}(x,y)\biggr)+R_{N}(t,x,y),$
and
$|R_{N}(t,x,y)|\leq C_{N}\,t^{N-n/2}$
for all $0<t\leq t_{0}$ and all $x,y\in M$ with $d_{g}(x,y)<r_{0}$. Moreover, there exist constants $c,C>0$ such that
(2.2) $0\leq p_{t}(x,y)\leq C\,t^{-n/2}\exp\Bigl(-\frac{d_{g}(x,y)^{2}}{ct}\Bigr),\qquad t\in(0,1],\ x,y\in M.$
Since $u_{0}$ is smooth and $u_{0}(x,x)=1$, there exists $C>0$ such that
(2.3) $|u_{0}(x,y)-1|\leq C\,d_{g}(x,y)$
whenever $d_{g}(x,y)<r_{0}$.
2.2. A smoothed Kuznecov sum
Define the heat-regularized Kuznecov sum associated with $\mu$ by
(2.4) $H_{\mu}(t):=\sum_{j=0}^{\infty}e^{-t\lambda_{j}^{2}}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2},\qquad t>0.$
We first relate $H_{\mu}(t)$ to the heat kernel.
Lemma 2.2.
For every $t>0$,
(2.5) $H_{\mu}(t)=\iint_{M\times M}p_{t}(x,y)\,d\mu(x)\,d\mu(y).$
Proof.
Let
$\alpha_{j}:=\int_{M}e_{j}\,d\mu,\qquad j=0,1,2,\dots.$
Recall that the heat kernel satisfies
$p_{t}(x,y)=\sum_{j=0}^{\infty}e^{-t\lambda_{j}^{2}}e_{j}(x)e_{j}(y).$
For each $J\in\mathbb{N}$ set
$p_{t}^{(J)}(x,y):=\sum_{j=0}^{J}e^{-t\lambda_{j}^{2}}e_{j}(x)e_{j}(y),\qquad S_{J}(t):=\sum_{j=0}^{J}e^{-t\lambda_{j}^{2}}|\alpha_{j}|^{2}.$
Since $e^{-t\lambda_{j}^{2}}\geq 0$ and $|\alpha_{j}|^{2}\geq 0$, the sequence $S_{J}(t)$ is nondecreasing in $J$.
For each fixed $t>0$, the series converges absolutely and uniformly on $M\times M$, since $\sup_{x,y}|e_{j}(x)e_{j}(y)|\leq\|e_{j}\|_{\infty}^{2}\lesssim\lambda_{j}^{n-1}$ by Hörmander [Hör68] and $\sum_{j\geq 0}e^{-t\lambda_{j}^{2}}\lambda_{j}^{n-1}<\infty$ by Weyl law. In particular,
$\sup_{x,y\in M}\bigl|p_{t}(x,y)-p_{t}^{(J)}(x,y)\bigr|\longrightarrow 0\qquad(J\to\infty),$
and $p_{t}$ is continuous and bounded on $M\times M$.
For each $J$ we have
$\begin{split}\iint_{M\times M}p_{t}^{(J)}(x,y)\,d\mu(x)\,d\mu(y)&=\iint_{M\times M}\sum_{j=0}^{J}e^{-t\lambda_{j}^{2}}e_{j}(x)e_{j}(y)\,d\mu(x)\,d\mu(y)\\ &=\sum_{j=0}^{J}e^{-t\lambda_{j}^{2}}\biggl(\int_{M}e_{j}(x)\,d\mu(x)\biggr)\biggl(\int_{M}e_{j}(y)\,d\mu(y)\biggr)\\ &=S_{J}(t),\end{split}$
where we have interchanged the finite sum and the integrals.
Since $p_{t}^{(J)}\to p_{t}$ uniformly and $\mu$ is finite, we have
$\lim_{J\to\infty}\iint_{M\times M}p_{t}^{(J)}(x,y)\,d\mu(x)\,d\mu(y)=\iint_{M\times M}p_{t}(x,y)\,d\mu(x)\,d\mu(y).$
Therefore $S_{J}(t)$ converges as $J\to\infty$, and
$\sum_{j=0}^{\infty}e^{-t\lambda_{j}^{2}}|\alpha_{j}|^{2}=\lim_{J\to\infty}S_{J}(t)=\iint_{M\times M}p_{t}(x,y)\,d\mu(x)\,d\mu(y),$
which is precisely (2.5). ∎
2.3. A Gaussian average
Using the distance distribution $F(r)$ introduced above, we consider the Gaussian average
$G(t):=\iint e^{-d_{g}(x,y)^{2}/(4t)}\,d\mu(x)\,d\mu(y).$
Let $R(x,y):=d_{g}(x,y)$ and let $\nu:=R_{*}(\mu\times\mu)$ be the pushforward of $\mu\times\mu$ under $R$. Then $\nu$ is a finite Borel measure on $0,\operatorname{diam}(M)]$ and
$F(r)=(\mu\times\mu)\bigl(\{(x,y):d_{g}(x,y)\leq r\}\bigr)=\nu([0,r]),$
so $F$ is the distribution function of $\nu$. In particular, for the nonnegative Borel function $\varphi(r):=e^{-r^{2}/(4t)}$ we have
$G(t)=\int_{M\times M}\varphi(R(x,y))\,d\mu(x)\,d\mu(y)=\int_{[0,\operatorname{diam}(M)]}\varphi(r)\,d\nu(r)=\int_{0}^{\operatorname{diam}(M)}e^{-r^{2}/(4t)}\,dF(r).$
The next lemma identifies the behavior of $G(t)$ as $t\to 0^{+}$.
Lemma 2.3.
Assume that $\mu$ satisfies (1.3) and (1.5). Then, as $t\to 0^{+}$,
(2.6) $G(t)=\gamma_{s}\,A_{\mu}\,t^{s/2}+o(t^{s/2}),$
where
(2.7) $\gamma_{s}:=s\,2^{\,s-1}\Gamma\Bigl(\frac{s}{2}\Bigr)\,\operatorname{vol}(B^{s}).$
In particular, there exists $C>0$ and $t_{1}>0$ such that
(2.8) $G(t)\leq C\,t^{s/2}\qquad(0<t\leq t_{1}).$
Proof.
Write, by our assumption on $\mu$,
$F(r)=\operatorname{vol}(B^{s})A_{\mu}r^{s}+\mathcal{E}(r),$
with $\mathcal{E}(r)=o(r^{s})$ as $r\to 0^{+}$. Fix $r_{0}>0$ as in (1.3). Split
$G(t)=\int_{0}^{r_{0}}e^{-r^{2}/(4t)}\,dF(r)+\int_{r_{0}}^{\operatorname{diam}(M)}e^{-r^{2}/(4t)}\,dF(r)=:G_{0}(t)+G_{1}(t).$
Since $F$ is bounded on $r_{0},\operatorname{diam}(M)]$, we have
$|G_{1}(t)|\leq F(\operatorname{diam}(M))\,e^{-r_{0}^{2}/(4t)}=o(t^{N})$
for every $N>0$ as $t\to 0^{+}$. In particular $G_{1}(t)=o(t^{s/2})$.
For $G_{0}(t)$, decompose
$G_{0}(t)=\operatorname{vol}(B^{s})A_{\mu}\int_{0}^{r_{0}}e^{-r^{2}/(4t)}\,d(r^{s})+\int_{0}^{r_{0}}e^{-r^{2}/(4t)}\,d\mathcal{E}(r)=:J_{1}(t)+J_{2}(t).$
For $J_{1}(t)$ we compute explicitly:
$J_{1}(t)=\operatorname{vol}(B^{s})A_{\mu}s\int_{0}^{r_{0}}r^{s-1}e^{-r^{2}/(4t)}\,dr=\operatorname{vol}(B^{s})A_{\mu}s\int_{0}^{\infty}r^{s-1}e^{-r^{2}/(4t)}\,dr+O(e^{-r_{0}^{2}/(4t)}).$
The error term is $o(t^{s/2})$ as above. Using the standard integral
$\int_{0}^{\infty}r^{s-1}e^{-r^{2}/(4t)}\,dr=2^{\,s-1}\Gamma\Bigl(\frac{s}{2}\Bigr)t^{s/2},$
we obtain
$J_{1}(t)=\gamma_{s}\,A_{\mu}\,t^{s/2}+o(t^{s/2})$
with $\gamma_{s}$ as in (2.7).
For $J_{2}(t)$ we integrate by parts on $0,r_{0}]$:
$J_{2}(t)=e^{-r^{2}/(4t)}\mathcal{E}(r)\Big|_{0}^{r_{0}}+\int_{0}^{r_{0}}\frac{r}{2t}e^{-r^{2}/(4t)}\mathcal{E}(r)\,dr=:B(t)+I(t).$
Since $F(0)=0$ and hence $\mathcal{E}(0)=0$, we have $B(t)=e^{-r_{0}^{2}/(4t)}\mathcal{E}(r_{0})=o(t^{s/2})$.
To handle $I(t)$, fix $\varepsilon>0$ and choose $\delta\in(0,r_{0})$ so small that $|\mathcal{E}(r)|\leq\varepsilon r^{s}$ for $0<r\leq\delta$. Split
$I(t)=\int_{0}^{\delta}\frac{r}{2t}e^{-r^{2}/(4t)}\mathcal{E}(r)\,dr+\int_{\delta}^{r_{0}}\frac{r}{2t}e^{-r^{2}/(4t)}\mathcal{E}(r)\,dr=:I_{1}(t)+I_{2}(t).$
On $(0,\delta]$ we have $|\mathcal{E}(r)|\leq\varepsilon r^{s}$, so
$|I_{1}(t)|\leq\frac{\varepsilon}{2t}\int_{0}^{\delta}r^{s+1}e^{-r^{2}/(4t)}\,dr\leq\frac{\varepsilon}{2t}\int_{0}^{\infty}r^{s+1}e^{-r^{2}/(4t)}\,dr=C_{1}\,\varepsilon\,t^{s/2}$
for some $C_{1}$ depending only on $s$.
On $\delta,r_{0}]$ we use the upper Ahlfors bound $F(r)\leq Cr^{s}$ to deduce that $|\mathcal{E}(r)|\leq C^{\prime}r^{s}$ on this interval. Thus
$|I_{2}(t)|\leq\frac{C^{\prime}}{2t}\int_{\delta}^{r_{0}}r^{s+1}e^{-r^{2}/(4t)}\,dr\leq C^{\prime\prime}t^{-1}e^{-\delta^{2}/(4t)},$
which is $o(t^{s/2})$ as $t\to 0^{+}$. Hence $I(t)=o(t^{s/2})$, and since $\varepsilon>0$ was arbitrary, we also have $J_{2}(t)=B(t)+I(t)=o(t^{s/2})$.
Combining the estimates for $J_{1}(t)$ and $J_{2}(t)$ gives (2.6). The upper bound (2.8) follows from (2.6) by taking $t$ sufficiently small and then enlarging the constant $C$ to cover all $t\in(0,t_{1}]$. ∎
2.4. Small-time asymptotics of the smoothed Kuznecov sum
We can now identify the leading behavior of $H_{\mu}(t)$ as $t\to 0^{+}$.
Lemma 2.4.
Assume that $\mu$ satisfies (1.3) and (1.5). Then
(2.9) $H_{\mu}(t)=\gamma_{n,s}\,A_{\mu}\,t^{-(n-s)/2}+o\bigl(t^{-(n-s)/2}\bigr)\qquad(t\to 0^{+}),$
where
(2.10) $\gamma_{n,s}:=(4\pi)^{-n/2}\,\gamma_{s}=(4\pi)^{-n/2}\,s\,2^{\,s-1}\Gamma\Bigl(\frac{s}{2}\Bigr)\,\operatorname{vol}(B^{s}).$
Proof.
By Lemma 2.2,
$H_{\mu}(t)=\iint_{M\times M}p_{t}(x,y)\,d\mu(x)\,d\mu(y).$
Fix $r_{0}>0$ as in Lemma 2.1 and split
$H_{\mu}(t)=\iint_{d_{g}(x,y)\leq r_{0}}p_{t}(x,y)\,d\mu(x)\,d\mu(y)+\iint_{d_{g}(x,y)>r_{0}}p_{t}(x,y)\,d\mu(x)\,d\mu(y)=:H_{0}(t)+H_{1}(t).$
By the Gaussian upper bound (2.2),
$0\leq H_{1}(t)\leq C\,t^{-n/2}\exp\Bigl(-\frac{r_{0}^{2}}{ct}\Bigr)\mu(M)^{2}=o\bigl(t^{-(n-s)/2}\bigr),$
since $t^{(n-s)/2}e^{-r_{0}^{2}/(ct)}\to 0$ as $t\to 0^{+}$.
For $H_{0}(t)$ we apply (2.1) with an integer $N\geq 1$ to be chosen momentarily. For $d_{g}(x,y)\leq r_{0}$ and $0<t\leq t_{0}$ we have
$p_{t}(x,y)=(4\pi t)^{-n/2}e^{-d_{g}(x,y)^{2}/(4t)}\biggl(\sum_{k=0}^{N-1}t^{k}u_{k}(x,y)\biggr)+R_{N}(t,x,y),$
with $|R_{N}(t,x,y)|\leq C_{N}t^{N-n/2}$. Thus
$H_{0}(t)=(4\pi t)^{-n/2}\sum_{k=0}^{N-1}t^{k}\iint_{d_{g}(x,y)\leq r_{0}}e^{-d_{g}(x,y)^{2}/(4t)}u_{k}(x,y)\,d\mu(x)\,d\mu(y)+E_{N}(t),$
where
$|E_{N}(t)|\leq C_{N}t^{N-n/2}\mu(M)^{2}.$
We first single out the $k=0$ term. Noting (2.3), we may set $\omega:=u_{0}-1$, and write
$\begin{split}&\iint_{d_{g}(x,y)\leq r_{0}}e^{-d_{g}(x,y)^{2}/(4t)}u_{0}(x,y)\,d\mu(x)\,d\mu(y)\\ &=\iint_{d_{g}(x,y)\leq r_{0}}e^{-d_{g}(x,y)^{2}/(4t)}\,d\mu(x)\,d\mu(y)+\iint_{d_{g}(x,y)\leq r_{0}}e^{-d_{g}(x,y)^{2}/(4t)}\omega(x,y)\,d\mu(x)\,d\mu(y).\end{split}$
The first integral differs from $G(t)$ by a tail integral:
$\begin{split}\iint_{d_{g}(x,y)\leq r_{0}}e^{-d_{g}(x,y)^{2}/(4t)}\,d\mu(x)\,d\mu(y)&=G(t)-\iint_{d_{g}(x,y)>r_{0}}e^{-d_{g}(x,y)^{2}/(4t)}\,d\mu(x)\,d\mu(y)\\ &=G(t)+T(t),\end{split}$
where
$|T(t)|\leq\mu(M)^{2}e^{-r_{0}^{2}/(4t)}=o(t^{s/2})$
as $t\to 0^{+}$.
For the second integral we use (2.3) and the Ahlfors bound via $F(r)$:
$\begin{split}\biggl|\iint_{d_{g}(x,y)\leq r_{0}}e^{-d_{g}(x,y)^{2}/(4t)}\omega(x,y)\,d\mu(x)\,d\mu(y)\biggr|&\leq C_{2}\iint_{d_{g}(x,y)\leq r_{0}}d_{g}(x,y)e^{-d_{g}(x,y)^{2}/(4t)}\,d\mu(x)\,d\mu(y)\\ &=C_{2}\int_{0}^{r_{0}}re^{-r^{2}/(4t)}\,dF(r)\\ &=:C_{2}J(t).\end{split}$
Similar to the proof of Lemma 2.3, noting $F(0)=0$ and (1.5), integration by parts gives
$|J(t)|\leq Cr_{0}^{s+1}e^{-r_{0}^{2}/(4t)}+C\int_{0}^{r_{0}}e^{-r^{2}/(4t)}r^{s}\,dr+\frac{C}{t}\int_{0}^{r_{0}}e^{-r^{2}/(4t)}r^{s+2}\,dr.$
Extending the integrals to $\infty$ and changing variables $r=\sqrt{t}\,u$ gives
$\int_{0}^{\infty}e^{-r^{2}/(4t)}r^{a}\,dr=t^{(a+1)/2}\int_{0}^{\infty}e^{-u^{2}/4}u^{a}\,du\lesssim t^{(a+1)/2},$
so the last two terms are $O(t^{(s+1)/2})$ and $t^{-1}O(t^{(s+3)/2})=O(t^{(s+1)/2})$, while the boundary term is $o(t^{(s+1)/2})$, and thus $J(t)\lesssim t^{(s+1)/2}$. Hence
$\iint_{d_{g}(x,y)\leq r_{0}}e^{-d_{g}(x,y)^{2}/(4t)}\omega(x,y)\,d\mu(x)\,d\mu(y)=O\bigl(t^{(s+1)/2}\bigr).$
Putting these estimates together, we obtain
$\iint_{d_{g}(x,y)\leq r_{0}}e^{-d_{g}(x,y)^{2}/(4t)}u_{0}(x,y)\,d\mu(x)\,d\mu(y)=G(t)+O\bigl(t^{(s+1)/2}\bigr).$
Therefore
$(4\pi t)^{-n/2}\iint_{d_{g}(x,y)\leq r_{0}}e^{-d_{g}(x,y)^{2}/(4t)}u_{0}(x,y)\,d\mu(x)\,d\mu(y)=(4\pi t)^{-n/2}G(t)+E_{0}(t),$
where
$|E_{0}(t)|\lesssim t^{-n/2}t^{(s+1)/2}=t^{1/2-(n-s)/2}=o\bigl(t^{-(n-s)/2}\bigr)$
as $t\to 0^{+}$.
Next, for each $k=1,\dots,N-1$, the smoothness of $u_{k}$ and the upper bound (2.8) yield
$\begin{split}&\biggl|(4\pi t)^{-n/2}t^{k}\iint_{d_{g}(x,y)\leq r_{0}}e^{-d_{g}(x,y)^{2}/(4t)}u_{k}(x,y)\,d\mu(x)\,d\mu(y)\biggr|\\ &\qquad\leq C_{k}t^{-n/2+k}\iint e^{-d_{g}(x,y)^{2}/(4t)}\,d\mu(x)\,d\mu(y)=C_{k}t^{-n/2+k}G(t)\lesssim t^{-n/2+k}t^{s/2}\\ &\qquad=O\bigl(t^{k-(n-s)/2}\bigr)=o\bigl(t^{-(n-s)/2}\bigr),\end{split}$
since $k\geq 1$.
Finally, for the remainder term we have
$|E_{N}(t)|\leq C_{N}t^{N-n/2}\mu(M)^{2}.$
Choose an integer $N>\frac{s}{2}$. Then
$t^{(n-s)/2}|E_{N}(t)|\lesssim t^{N-s/2}\longrightarrow 0\qquad(t\to 0^{+}),$
so $E_{N}(t)=o\bigl(t^{-(n-s)/2}\bigr)$.
Collecting the contributions from $k=0,\dots,N-1$, the remainder $E_{N}$, and $H_{1}(t)$, we conclude that
$H_{\mu}(t)=(4\pi t)^{-n/2}G(t)+o\bigl(t^{-(n-s)/2}\bigr)\qquad(t\to 0^{+}).$
Using Lemma 2.3, which gives $G(t)=\gamma_{s}A_{\mu}t^{s/2}+o(t^{s/2})$, we obtain
$H_{\mu}(t)=(4\pi)^{-n/2}\gamma_{s}A_{\mu}t^{-(n-s)/2}+o\bigl(t^{-(n-s)/2}\bigr),$
which is exactly (2.9) with $\gamma_{n,s}$ given by (2.10). ∎
3. A Tauberian theorem and the proof of Theorem 1.3
We now view the smoothed Kuznecov sum $H_{\mu}(t)$ as a Laplace–Stieltjes transform of a spectral counting function and apply a Tauberian theorem to deduce the asymptotic behavior of $N_{\mu}(\lambda)$.
3.1. Laplace–Stieltjes representation
Define a nondecreasing, right-continuous function $\mathcal{N}:[0,\infty)\to[0,\infty)$ by
$\mathcal{N}(S):=\sum_{\lambda_{j}^{2}\leq S}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2},\qquad S\geq 0.$
Thus $\mathcal{N}$ puts mass $|\int_{M}e_{j}\,d\mu|^{2}$ at the point $S=\lambda_{j}^{2}$. By definition (1.1),
(3.1) $\mathcal{N}(\lambda^{2})=N_{\mu}(\lambda)\qquad(\lambda\geq 0).$
Using Lemma 2.2 and the spectral expansion of $p_{t}$, we have
(3.2) $H_{\mu}(t)=\sum_{j=0}^{\infty}e^{-t\lambda_{j}^{2}}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}=\int_{[0,\infty)}e^{-tS}\,d\mathcal{N}(S),\qquad t>0.$
In particular, $H_{\mu}(t)$ is the Laplace–Stieltjes transform of $\mathcal{N}$.
3.2. A Tauberian theorem for Laplace–Stieltjes transforms
We now recall a standard Tauberian theorem which connects the singular behavior of a Laplace–Stieltjes transform at the origin to the growth of the underlying measure. This is a special case of Karamata’s theorem. See, for example, [BGT87, Theorem 1.7.1].
Theorem 3.1 (Karamata’s Tauberian Theorem).
Let $\mathcal{N}:[0,\infty)\to[0,\infty)$ be nondecreasing and right-continuous, and let
$\mathcal{F}(t):=\int_{[0,\infty)}e^{-tS}\,d\mathcal{N}(S)$
be finite for all $t>0$. Suppose that for some constants $C>0$ and $\beta>0$,
$\mathcal{F}(t)\sim C\,t^{-\beta}\qquad(t\to 0^{+}).$
Then
$\mathcal{N}(S)\sim\frac{C}{\Gamma(\beta+1)}\,S^{\beta}\qquad(S\to\infty).$
Theorem 3.1 is the Tauberian half of the usual “Abelian $\Longleftrightarrow$ Tauberian” correspondence for Laplace–Stieltjes transforms with index $\beta>0$: roughly speaking, asymptotics of $\mathcal{N}(S)$ as $S\to\infty$ and asymptotics of its Laplace–Stieltjes transform $\mathcal{F}(t)$ as $t\to 0^{+}$ determine each other. A proof in this general framework can be found in [BGT87, §1.7].
3.3. Proof of Theorem 1.3
We now combine Lemma 2.4 with Theorem 3.1. Recall that Lemma 2.4 shows that
$H_{\mu}(t)=\gamma_{n,s}\,A_{\mu}\,t^{-(n-s)/2}+o\bigl(t^{-(n-s)/2}\bigr)\qquad(t\to 0^{+}),$
with $\gamma_{n,s}$ given by (2.10). In the notation of Theorem 3.1 we have
$\mathcal{F}(t)=H_{\mu}(t),\quad C=\gamma_{n,s}A_{\mu},\quad\beta=\frac{n-s}{2}.$
Therefore
$\mathcal{N}(S)\sim\frac{\gamma_{n,s}A_{\mu}}{\Gamma\bigl(\frac{n-s}{2}+1\bigr)}\,S^{(n-s)/2}\qquad(S\to\infty).$
Using (3.1) with $S=\lambda^{2}$, we obtain
$N_{\mu}(\lambda)=\mathcal{N}(\lambda^{2})\sim\frac{\gamma_{n,s}A_{\mu}}{\Gamma\bigl(\tfrac{n-s}{2}+1\bigr)}\,\lambda^{n-s}\qquad(\lambda\to\infty).$
In particular, (1.7) holds with
$C_{n,s}=\frac{\gamma_{n,s}}{\Gamma\bigl(\tfrac{n-s}{2}+1\bigr)}.$
It remains to simplify this constant. Recall from (2.10) and (2.7) that
$\gamma_{n,s}=(4\pi)^{-n/2}\,s\,2^{\,s-1}\Gamma\Bigl(\frac{s}{2}\Bigr)\,\operatorname{vol}(B^{s}),$
and that
$\operatorname{vol}(B^{m})=\frac{\pi^{m/2}}{\Gamma\bigl(\frac{m}{2}+1\bigr)},\qquad m>0.$
Since $\Gamma\bigl(\frac{s}{2}+1\bigr)=\frac{s}{2}\Gamma\bigl(\frac{s}{2}\bigr)$, we may write
$\operatorname{vol}(B^{s})=\frac{\pi^{s/2}}{\Gamma\bigl(\frac{s}{2}+1\bigr)}=\frac{2}{s}\,\pi^{s/2}\,\Gamma\Bigl(\frac{s}{2}\Bigr)^{-1}.$
Substituting this into the formula for $\gamma_{n,s}$ gives
$\gamma_{n,s}=(4\pi)^{-n/2}\,s\,2^{\,s-1}\Gamma\Bigl(\frac{s}{2}\Bigr)\cdot\frac{2}{s}\,\pi^{s/2}\,\Gamma\Bigl(\frac{s}{2}\Bigr)^{-1}=(4\pi)^{-n/2}\,2^{\,s}\,\pi^{s/2}.$
Hence
$C_{n,s}=\frac{(4\pi)^{-n/2}2^{\,s}\pi^{s/2}}{\Gamma\bigl(\frac{n-s}{2}+1\bigr)}=2^{-(n-s)}\pi^{-(n-s)/2}\frac{1}{\Gamma\bigl(\frac{n-s}{2}+1\bigr)}.$
On the other hand, taking m=nsm=n-s in the volume formula,
$\operatorname{vol}(B^{n-s})=\frac{\pi^{(n-s)/2}}{\Gamma\bigl(\frac{n-s}{2}+1\bigr)},$
so
$(2\pi)^{-(n-s)}\,\operatorname{vol}(B^{n-s})=2^{-(n-s)}\pi^{-(n-s)}\cdot\frac{\pi^{(n-s)/2}}{\Gamma\bigl(\frac{n-s}{2}+1\bigr)}=2^{-(n-s)}\pi^{-(n-s)/2}\frac{1}{\Gamma\bigl(\frac{n-s}{2}+1\bigr)}=C_{n,s}.$
Thus
$C_{n,s}=(2\pi)^{-(n-s)}\,\operatorname{vol}(B^{\,n-s}),$
which is exactly the constant stated in Theorem 1.3. This completes the proof.
4. formulae of Hare–Roginskaya and the proof of Theorem 1.5
In this section we explain how the energy identities of Hare and Roginskaya [HR03] imply the polynomial upper bounds stated in Theorem 1.5.
Recall that for $0<u<n$ we defined the Riesz energy of a finite Borel measure $\mu$ on $M$ by
$I_{u}(\mu)=\iint d_{g}(x,y)^{-u}\,d\mu(x)\,d\mu(y),$
where $d_{g}(x,y)$ denotes the Riemannian distance on $(M,g)$, see (1.8). Hare and Roginskaya relate $I_{u}(\mu)$ to a weighted spectral sum via the kernel of a suitable elliptic operator. For our purposes, we only need the following consequence of their results.
Proposition 4.1 ([HR03]).
Let $(M,g)$ be a compact Riemannian manifold of dimension $n\geq 2$. For each $u\in(0,n)$ there exists a constant $c_{n,u}\geq 1$, depending only on $n$, $u$ and $(M,g)$, such that for every finite Borel measure $\mu$ on $M$,
(4.1) $c_{n,u}^{-1}I_{u}(\mu)\leq\sum_{j=0}^{\infty}(1+\lambda_{j}^{2})^{-(n-u)/2}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}\leq c_{n,u}\,I_{u}(\mu).$
Remark 4.2.
Estimates of the form (4.1) are proved in [HR03] by expressing suitable powers of the Laplace–Beltrami operator as pseudodifferential operators and identifying their kernels with Riesz-type potentials. In particular, $I_{u}(\mu)$ is comparable, up to multiplicative constants depending only on $(M,g,n,u)$, to the quadratic form associated with the operator $(I-\Delta_{g})^{-(n-u)/2}$ evaluated on $\mu$. See [HR03] for details.
Given Proposition 4.1, the proof of Theorem 1.5 is straightforward.
Proof of Theorem 1.5
Let $u\in(0,n)$ and let $\mu$ be a finite Borel measure with $I_{u}(\mu)<\infty$. By the upper bound in (4.1),
(4.2) $\sum_{j=0}^{\infty}(1+\lambda_{j}^{2})^{-(n-u)/2}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}\leq c_{n,u}\,I_{u}(\mu).$
Now fix $\lambda\geq 1$. Since the weight $(1+\lambda_{j}^{2})^{-(n-u)/2}$ is decreasing in $\lambda_{j}$, we have
$(1+\lambda_{j}^{2})^{-(n-u)/2}\geq(1+\lambda^{2})^{-(n-u)/2}\qquad\text{whenever }\lambda_{j}\leq\lambda,$
and hence
$\sum_{\lambda_{j}\leq\lambda}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}\leq(1+\lambda^{2})^{\frac{n-u}{2}}\sum_{j=0}^{\infty}(1+\lambda_{j}^{2})^{-(n-u)/2}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}.$
Combining this with (4.2) gives
$\sum_{\lambda_{j}\leq\lambda}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}\leq C_{1}(M,g,u,\mu)\,(1+\lambda^{2})^{\frac{n-u}{2}}\leq C_{1}(M,g,u,\mu)\,\lambda^{\,n-u}$
for all $\lambda\geq 1$, where we may take
$C_{1}(M,g,u,\mu):=2^{\frac{n-u}{2}}c_{n,u}\,I_{u}(\mu).$
This proves the first assertion of Theorem 1.5.
To obtain the “in particular” statement, it remains to note that if $\mu$ is $s$-Ahlfors regular for some $s\in(0,n)$, then $I_{u}(\mu)<\infty$ for every $u<s$. This is a standard estimate, but we include a brief argument in our setting for completeness.
Lemma 4.3.
Let $\mu$ be an $s$-Ahlfors regular measure on $M$ with $s\in(0,n)$, and let $u\in(0,s)$. Then $I_{u}(\mu)<\infty$.
Proof.
Recall
$F(r)=(\mu\times\mu)\bigl(\{(x,y):d_{g}(x,y)\leq r\}\bigr)=\int_{M}\mu(B(x,r))\,d\mu(x).$
Using the layer-cake representation, we have
$I_{u}(\mu)=\int_{0}^{\operatorname{diam}(M)}ur^{-u-1}F(r)\,dr+\mu(M)^{2}\,\operatorname{diam}(M)^{-u},$
Split the integral at $r_{0}$:
$I_{u}(\mu)=u\int_{0}^{r_{0}}r^{-u-1}F(r)\,dr+u\int_{r_{0}}^{\operatorname{diam}(M)}r^{-u-1}F(r)\,dr+\mu(M)^{2}\,\operatorname{diam}(M)^{-u}.$
On $(0,r_{0}]$,
$u\int_{0}^{r_{0}}r^{-u-1}F(r)\,dr\leq uC\int_{0}^{r_{0}}r^{s-u-1}\,dr=\frac{uC}{s-u}\,r_{0}^{\,s-u}<\infty,$
since su>0s-u>0. On $r_{0},\operatorname{diam}(M)]$,
$u\int_{r_{0}}^{\operatorname{diam}(M)}r^{-u-1}F(r)\,dr\leq u\mu(M)^{2}\int_{r_{0}}^{\operatorname{diam}(M)}r^{-u-1}\,dr<\infty.$
Thus $I_{u}(\mu)<\infty$ for every $u<s$, as claimed. ∎
Applying Lemma 4.3 with any fixed $u<s$ shows that the first part of Theorem 1.5 applies to $\mu$ and yields a constant $C_{2}=C_{2}(M,g,s,u,\mu)$ such that
$\sum_{\lambda_{j}\leq\lambda}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}\leq C_{2}\,\lambda^{\,n-u}\qquad(\lambda\geq 1),$
without any density assumption. This completes the proof of Theorem 1.5.
Remark 4.4.
Theorem 1.5 shows that for each fixed $u<s$ one can control the growth of the Kuznecov sum by $\lambda^{n-u}$ under the mild hypothesis $I_{u}(\mu)<\infty$, and in particular for $s$-Ahlfors regular measures. Our main theorem, Theorem 1.3, goes further by identifying the sharp exponent nsn-s and the exact leading constant in the asymptotic (1.7), under the additional assumption of the averaged $s$-density condition. As discussed in the introduction, the methods of [HR03] treat $u$ as a fixed parameter and do not provide the uniform control in $u$ near $s$ that is needed to pass to the endpoint $u=s$. This is why we turn to a more refined analysis based on the heat kernel and Tauberian theory.
5. Sharpness of Theorem 1.3
In this section, we discuss the sharpness of Theorem 1.3. We explain the necessity of the two structural hypotheses, namely $s$-Ahlfors regularity and the averaged $s$-density condition (1.5), for obtaining the exponent nsn-s and a genuine one-term asymptotic. We also show that the remainder $o(\lambda^{n-s})$ in (1.7) is essentially optimal at this level of generality: in general it cannot be improved uniformly to a power-saving error term over the class of measures under consideration.
5.1. Necessity of the Ahlfors regularity assumption
It is natural to ask whether the exponent nsn-s in Theorem 1.3 (or in the upper bound of Theorem 1.5) could be expressed solely in terms of the Hausdorff dimension of $\mu$, without any Ahlfors regularity hypothesis. This is not possible in general.
Let $H_{1},H_{2}\subset M$ be disjoint embedded submanifolds with dimensions $s_{1}<s_{2}$, and let $\mu_{i}=dV_{H_{i}}$ and $\mu=\mu_{1}+\mu_{2}$. Then
$\dim_{H}(\mu)=\max(s_{1},s_{2})=s_{2},$
but the growth of $N_{\mu}(\lambda)$ is governed by the smaller dimension $s_{1}$.
To see this, write
$a_{j}:=\int_{H_{1}}e_{j}\,dV_{H_{1}},\qquad b_{j}:=\int_{H_{2}}e_{j}\,dV_{H_{2}}.$
Then
$N_{\mu}(\lambda)=\sum_{\lambda_{j}\leq\lambda}|a_{j}+b_{j}|^{2}=\sum_{\lambda_{j}\leq\lambda}|a_{j}|^{2}+\sum_{\lambda_{j}\leq\lambda}|b_{j}|^{2}+2\Re\sum_{\lambda_{j}\leq\lambda}a_{j}\overline{b_{j}}.$
By Cauchy–Schwarz,
$\Bigl|\sum_{\lambda_{j}\leq\lambda}a_{j}\overline{b_{j}}\Bigr|\leq\Bigl(\sum_{\lambda_{j}\leq\lambda}|a_{j}|^{2}\Bigr)^{1/2}\Bigl(\sum_{\lambda_{j}\leq\lambda}|b_{j}|^{2}\Bigr)^{1/2}.$
Hence we have the robust comparison
(5.1) $\bigl|N_{\mu}(\lambda)-N_{\mu_{1}}(\lambda)-N_{\mu_{2}}(\lambda)\bigr|\leq 2\,\bigl(N_{\mu_{1}}(\lambda)N_{\mu_{2}}(\lambda)\bigr)^{1/2}.$
By the smooth submanifold Kuznecov formulae,
$N_{\mu_{i}}(\lambda)\sim C_{n,s_{i}}\,\operatorname{vol}(H_{i})\,\lambda^{n-s_{i}}\qquad(i=1,2),$
so $N_{\mu_{2}}(\lambda)=o(N_{\mu_{1}}(\lambda))$ and the right-hand side of (5.1) is
$O\Bigl(\lambda^{\frac{n-s_{1}}{2}}\,\lambda^{\frac{n-s_{2}}{2}}\Bigr)=O\bigl(\lambda^{\,n-\frac{s_{1}+s_{2}}{2}}\bigr)=o\bigl(\lambda^{n-s_{1}}\bigr).$
Therefore $N_{\mu}(\lambda)=N_{\mu_{1}}(\lambda)+o(\lambda^{n-s_{1}})$, and in particular
$N_{\mu}(\lambda)\sim C_{n,s_{1}}\,\operatorname{vol}(H_{1})\,\lambda^{n-s_{1}}\qquad(\lambda\to\infty),$
even though $\dim_{H}(\mu)=s_{2}$. Thus, no exponent depending only on $\dim_{H}(\mu)$ can govern $N_{\mu}(\lambda)$ for arbitrary finite Borel measures.
5.2. Necessity of the averaged $s$-density condition
The averaged $s$-density hypothesis in Theorem 1.3 is not merely a technical convenience: without a stabilization of the small-scale behavior of $\mu$, one cannot in general expect the normalized quantities $N_{\mu}(\lambda)/\lambda^{n-s}$ to converge.
Recall the distance distribution
$F(r):=(\mu\times\mu)\bigl(\{(x,y):d_{g}(x,y)\leq r\}\bigr),\qquad r\geq 0,$
and the Gaussian average
$G(t):=\iint e^{-d_{g}(x,y)^{2}/(4t)}\,d\mu(x)\,d\mu(y)=\int_{0}^{\operatorname{diam}(M)}e^{-r^{2}/(4t)}\,dF(r).$
Define the normalized coefficient
$a(r):=\frac{F(r)}{r^{s}}\qquad(r>0).$
If $\mu$ is $s$-Ahlfors regular then $a(r)$ is bounded above and below by positive constants for $0<r\leq r_{0}$.
Since $F(0)=0$, integration by parts gives, modulo a boundary term that is exponentially decaying in $1/t$, we have
(5.2) $G(t)=\frac{1}{2t}\int_{0}^{\operatorname{diam}(M)}re^{-r^{2}/(4t)}\,F(r)\,dr.$
Changing variables $r=\sqrt{t}\,u$ yields
(5.3) $\frac{G(t)}{t^{s/2}}=\int_{0}^{\operatorname{diam}(M)/\sqrt{t}}w_{s}(u)\,a(\sqrt{t}\,u)\,du,\qquad w_{s}(u):=\frac{1}{2}\,u^{s+1}e^{-u^{2}/4}.$
The weight $w_{s}$ is positive and integrable on $(0,\infty)$, with
$\kappa_{s}:=\int_{0}^{\infty}w_{s}(u)\,du=\frac{1}{2}\int_{0}^{\infty}u^{s+1}e^{-u^{2}/4}\,du=s\,2^{s-1}\Gamma\Bigl(\frac{s}{2}\Bigr).$
Let
$a_{-}:=\liminf_{r\to 0^{+}}a(r),\qquad a_{+}:=\limsup_{r\to 0^{+}}a(r).$
Using (5.3), the boundedness of $a$, and $\int_{0}^{\infty}w_{s}<\infty$, one obtains the general bounds
(5.4) $\kappa_{s}\,a_{-}\leq\liminf_{t\to 0^{+}}\frac{G(t)}{t^{s/2}}\leq\limsup_{t\to 0^{+}}\frac{G(t)}{t^{s/2}}\leq\kappa_{s}\,a_{+}.$
In general (5.4) may be strict: the Gaussian average is a weighted average of the values of $a(r)$ on the scale $r\approx\sqrt{t}$, so thin oscillations in $a$ can be smoothed.
However, in the model constructions relevant for sharpness one can arrange that different limit points of $a(r)$ persist across long blocks of scales. In that setting the endpoints in (5.4) are realized along suitable subsequences.
Lemma 5.1 (Realizing $\liminf/\limsup$ from symmetric scale blocks).
Assume $a$ is bounded on $(0,\operatorname{diam}(M)]$, and let
$a_{-}:=\liminf_{r\to 0^{+}}a(r),\qquad a_{+}:=\limsup_{r\to 0^{+}}a(r).$
  1. (1)
    If there exist $r_{m}\to 0^{+}$ and $U_{m}\to\infty$ such that
    $\sup_{\,r_{m}/U_{m}\leq r\leq U_{m}r_{m}}\,|a(r)-a_{-}|\longrightarrow 0,$
    then with $t_{m}:=r_{m}^{2}$ one has
    $\frac{G(t_{m})}{t_{m}^{s/2}}\longrightarrow\kappa_{s}\,a_{-}.$
  2. (2)
    If there exist $\rho_{m}\to 0^{+}$ and $V_{m}\to\infty$ such that
    $\sup_{\,\rho_{m}/V_{m}\leq r\leq V_{m}\rho_{m}}\,|a(r)-a_{+}|\longrightarrow 0,$
    then with $\tau_{m}:=\rho_{m}^{2}$ one has
    $\frac{G(\tau_{m})}{\tau_{m}^{s/2}}\longrightarrow\kappa_{s}\,a_{+}.$
Proof.
We prove (1); (2) is identical. Put $t_{m}=r_{m}^{2}$ and use (5.3):
$\frac{G(t_{m})}{t_{m}^{s/2}}=\int_{0}^{\operatorname{diam}(M)/r_{m}}w_{s}(u)\,a(r_{m}u)\,du.$
Split the integral into three pieces:
$\int_{0}^{\operatorname{diam}(M)/r_{m}}=\int_{0}^{1/U_{m}}+\int_{1/U_{m}}^{U_{m}}+\int_{U_{m}}^{\operatorname{diam}(M)/r_{m}}.$
On $u\in[1/U_{m},U_{m}]$ we have $r_{m}u\in[r_{m}/U_{m},\,U_{m}r_{m}]$, hence $a(r_{m}u)\to a_{-}$ uniformly, so
$\int_{1/U_{m}}^{U_{m}}w_{s}(u)\,a(r_{m}u)\,du=a_{-}\int_{1/U_{m}}^{U_{m}}w_{s}(u)\,du+o(1).$
Since $a$ is bounded, say $|a|\leq K$ on $(0,r_{0}]$, the tails satisfy
$\left|\int_{0}^{1/U_{m}}w_{s}(u)\,a(r_{m}u)\,du\right|\leq K\int_{0}^{1/U_{m}}w_{s}(u)\,du\longrightarrow 0,$
because $w_{s}(u)\approx u^{s+1}$ near 0, and
$\left|\int_{U_{m}}^{\operatorname{diam}(M)/r_{m}}w_{s}(u)\,a(r_{m}u)\,du\right|\leq K\int_{U_{m}}^{\infty}w_{s}(u)\,du\longrightarrow 0,$
because $w_{s}\in L^{1}(0,\infty)$ and $U_{m}\to\infty$. Finally,
$\int_{1/U_{m}}^{U_{m}}w_{s}(u)\,du\longrightarrow\int_{0}^{\infty}w_{s}(u)\,du=\kappa_{s},$
so $G(t_{m})/t_{m}^{s/2}\to\kappa_{s}a_{-}$. ∎
Lemma 5.1 shows that if $a(r)$ has two distinct limit points $a_{-}<a_{+}$ and one can arrange two sequences of logarithmic blocks on which $a(r)$ stays uniformly close to $a_{-}$ and $a_{+}$, respectively, then $G(t)/t^{s/2}$ has two distinct subsequential limits $\kappa_{s}a_{-}$ and $\kappa_{s}a_{+}$, and hence does not converge. By the near-diagonal reduction in Section 2, this forces $t^{(n-s)/2}H_{\mu}(t)$, and therefore $N_{\mu}(\lambda)/\lambda^{n-s}$, to fail to converge as well. This explains why some stabilization of $F(r)/r^{s}$ is needed for a genuine one-term asymptotic.
5.3. Optimality of the remainder
In the smooth submanifold case, Zelditch’s Kuznecov formulae (1.2) come with an $O(\lambda^{n-k-1})$ remainder term. Moreover, for generic metrics this remainder can be improved to $o(\lambda^{n-k-1})$ by a recent result of Kaloshin, Wyman and the author [KWX25]. In contrast, the asymptotic formula in Theorem 1.3 is stated with an $o(\lambda^{n-s})$ remainder. It is natural to ask whether, under the assumptions of $s$-Ahlfors regularity and the existence of an averaged $s$-density constant, one can improve this to a uniform power-saving error term of the form
(5.5) $N_{\mu}(\lambda)=C_{n,s}A_{\mu}\,\lambda^{n-s}+O(\lambda^{n-s-\delta})\qquad(\lambda\to\infty),$
for some $\delta>0$ depending only on $(M,g,n,s)$ and valid for all measures $\mu$ in our class. We now explain why such a uniform improvement is not available in general.
Step 1: from $N_{\mu}$ to the Laplace–Stieltjes transform $H_{\mu}$. Write $\mathcal{N}(S):=N_{\mu}(\sqrt{S})$ as in (3.1), so that
$H_{\mu}(t)=\int_{[0,\infty)}e^{-tS}\,d\mathcal{N}(S).$
If (5.5) holds for some $\delta>0$, then
$\mathcal{N}(S)=C_{n,s}A_{\mu}\,S^{\frac{n-s}{2}}+O\Bigl(S^{\frac{n-s-\delta}{2}}\Bigr),$
and a standard Abelian (integration-by-parts) computation yields
(5.6) $H_{\mu}(t)=C_{n,s}A_{\mu}\,\Gamma\Bigl(\frac{n-s}{2}+1\Bigr)\,t^{-\frac{n-s}{2}}+O\Bigl(t^{-\frac{n-s-\delta}{2}}\Bigr)\qquad(t\to 0^{+}).$
Step 2: from $H_{\mu}$ to the Gaussian average $G(t)$. Recall
$G(t):=\iint e^{-d_{g}(x,y)^{2}/(4t)}\,d\mu(x)\,d\mu(y).$
Inspecting the proof of Lemma 2.4 (using the Lipschitz bound (2.3), the estimate $J(t)\lesssim t^{(s+1)/2}$, and choosing an integer $N\geq\lceil(s+1)/2\rceil$ in the local heat-kernel parametrix) yields the quantitative refinement
(5.7) $H_{\mu}(t)=(4\pi t)^{-n/2}G(t)+O\!\left(t^{-\frac{n-s}{2}+\frac{1}{2}}\right)\qquad(t\to 0^{+}).$
Combining (5.6) and (5.7), and using that $C_{n,s}=\gamma_{n,s}/\Gamma(\frac{n-s}{2}+1)$ and $\gamma_{n,s}=(4\pi)^{-n/2}\gamma_{s}$, one obtains (for $\delta\in(0,1)$)
(5.8) $\frac{G(t)}{t^{s/2}}=\gamma_{s}A_{\mu}+O(t^{\delta/2})\qquad(t\to 0^{+}),$
where $\gamma_{s}$ is as in Lemma 2.3. Thus, a uniform power saving in (5.5) would force a uniform power saving in $t$ for $G(t)/t^{s/2}$.
Step 3: from $G(t)$ to $F(r)$. Recall $F(r):=(\mu\times\mu)\{(x,y):d_{g}(x,y)\leq r\}$ and set
$q(r):=\frac{F(r)}{\operatorname{vol}(B^{s})\,A_{\mu}\,r^{s}}.$
Then by definition, $q(r)\to 1$ as $r\to 0^{+}$. Integrating by parts in $G(t)=\int_{0}^{\operatorname{diam}(M)}e^{-r^{2}/(4t)}\,dF(r)$ and changing variables $r=\sqrt{t}\,u$ yields
(5.9) $\frac{G(t)}{\operatorname{vol}(B^{s})\,A_{\mu}\,t^{s/2}}=\int_{0}^{\operatorname{diam}(M)/\sqrt{t}}w_{s}(u)\,q(\sqrt{t}\,u)\,du+O(e^{-c/t}),$
where
$w_{s}(u):=\frac{1}{2}\,u^{s+1}e^{-u^{2}/4}$
is strictly positive and integrable on $(0,\infty)$.
Lemma 5.2.
Assume $q(r)\to 1$ as $r\to 0^{+}$ and $\|q\|_{L^{\infty}(0,\operatorname{diam}(M)]}\leq C_{0}$ for some $C_{0}\geq 1$. Fix $\eta\in(0,1)$ and choose $U=U(\eta)\geq 2$ so that
(5.10) $\int_{U}^{\infty}w_{s}(u)\,du\leq\frac{\eta}{8C_{0}}\int_{1}^{U}w_{s}(u)\,du.$
Fix $r>0$, put $t=r^{2}$, and assume that $\text{for all }\rho\in[r,Ur].$
$q(\rho)\leq 1-\eta.$
Then for all sufficiently small $r$ one has
$\left|\frac{G(t)}{t^{s/2}}-\gamma_{s}A_{\mu}\right|\geq c_{s}\,A_{\mu}\,\eta\int_{1}^{U}w_{s}(u)\,du,$
where $c_{s}>0$ depends only on $s$.
Proof.
Let $\kappa_{s}:=\int_{0}^{\infty}w_{s}(u)\,du$, so $\gamma_{s}=\operatorname{vol}(B^{s})\kappa_{s}$. Using (5.9) with $t=r^{2}$ and absorbing the exponentially small term, we get
$\frac{G(r^{2})}{\operatorname{vol}(B^{s})\,A_{\mu}\,r^{s}}-\kappa_{s}=\int_{0}^{\infty}w_{s}(u)\bigl(q(ru)-1\bigr)\,du+o(1)\qquad(r\to 0^{+}).$
Split the integral into $u\in(0,1)$, $u\in[1,U]$, and $u\in[U,\infty)$.
On $[1,U]$ we have $ru\in[r,Ur]$, so q(ru)1q(ru)-1 has fixed sign and $|q(ru)-1|\geq\eta$. Hence
$\left|\int_{1}^{U}w_{s}(u)\bigl(q(ru)-1\bigr)\,du\right|\geq\eta\int_{1}^{U}w_{s}(u)\,du.$
On $U,\infty)$ we use $|q-1|\leq C_{0}+1\leq 2C_{0}$ and (5.10):
$\left|\int_{U}^{\infty}w_{s}(u)\bigl(q(ru)-1\bigr)\,du\right|\leq 2C_{0}\int_{U}^{\infty}w_{s}(u)\,du\leq\frac{\eta}{4}\int_{1}^{U}w_{s}(u)\,du.$
Finally, since $q(\rho)\to 1$ as $\rho\to 0^{+}$ and $w_{s}\in L^{1}(0,1)$, dominated convergence gives
$\int_{0}^{1}w_{s}(u)\bigl(q(ru)-1\bigr)\,du\longrightarrow 0\qquad(r\to 0^{+}).$
Therefore for all sufficiently small $r$,
$\left|\int_{0}^{\infty}w_{s}(u)\bigl(q(ru)-1\bigr)\,du\right|\geq\frac{\eta}{2}\int_{1}^{U}w_{s}(u)\,du.$
Multiplying by $\operatorname{vol}(B^{s})A_{\mu}$ and using $\gamma_{s}=\operatorname{vol}(B^{s})\kappa_{s}$ yields the claim. ∎
Step 4: constructing arbitrarily slow convergence. We now construct an explicit example (for a fixed $\delta>0$) for which no uniform remainder $O(\lambda^{n-s-\delta})$ can hold. For simplicity, we take $s$ to be an integer, and
$M=\mathbb{T}^{n},\qquad H=\mathbb{T}^{s}\subset\mathbb{T}^{n}$
with the flat metric, so that $dV_{H}$ is Lebesgue measure on $\mathbb{T}^{s}$ and, for all $\rho\in(0,1/10]$, the intrinsic ball $B_{H}(x,\rho)$ is a translate of the Euclidean ball. We remark that similar constructions can be made for a general manifold–submanifold pair $(M,H)$. Let
$K_{\rho}(y):=\mathbf{1}_{B_{H}(0,\rho)}(y),\qquad\mathcal{A}_{\rho}f(x):=\frac{1}{\operatorname{vol}(B^{s})\rho^{s}}\int_{\mathbb{T}^{s}}f(x-y)K_{\rho}(y)\,dV_{H}(y).$
Then whenever $d\mu=w\,dV_{H}$ we have
$\frac{F(\rho)}{\operatorname{vol}(B^{s})\rho^{s}}=\int_{H}w(x)\,\mathcal{A}_{\rho}w(x)\,dV_{H}(x).$
Choose frequencies $k_{m}\in\mathbb{Z}^{s}\setminus\{0\}$ such that $|k_{m}|$ is strictly increasing and $k_{m}\neq\pm k_{m^{\prime}}$ for $m\neq m^{\prime}$, and define
$\phi_{m}(x):=\cos(2\pi k_{m}\cdot x),\qquad\|\phi_{m}\|_{L^{\infty}(H)}\leq 1,\qquad\int_{H}\phi_{m}\,dV_{H}=0,\qquad\int_{H}\phi_{m}^{2}\,dV_{H}=\tfrac{1}{2}\,\operatorname{vol}(H).$
For $\rho\in(0,1/10]$ the averaging convolution operator $\mathcal{A}_{\rho}$ acts diagonally on Fourier modes: there exists a continuous function $m_{s}:[0,\infty)\to\mathbb{R}$ with
$m_{s}(0)=1,\qquad\lim_{\tau\to\infty}m_{s}(\tau)=0,\qquad|m_{s}(\tau)|\leq 1,$
such that
$\mathcal{A}_{\rho}\phi_{m}=m_{s}(2\pi|k_{m}|\rho)\,\phi_{m}.$
Fix amplitudes $a_{m}>0$ such that
(5.11) $\sum_{m\geq 1}a_{m}\leq\tfrac{1}{4},\qquad\text{and hence}\qquad\sum_{m\geq 1}a_{m}^{2}<\infty.$
Define
$w(x):=1+\sum_{m\geq 1}a_{m}\,\phi_{m}(x),\qquad\text{and}\qquad d\mu:=w\,dV_{H}.$
Then $\tfrac{3}{4}\leq w\leq\tfrac{5}{4}$, so $\mu$ is $s$-Ahlfors regular. Moreover, since $\mu$ is absolutely continuous with respect to $dV_{H}$ with bounded density $w$, $\mu$ admits a pointwise $s$-density $\theta=w$. Hence, by Remark 1.4,
$A_{\mu}=\int_{H}\theta\,d\mu=\int_{H}w^{2}\,dV_{H}.$
Using orthogonality of distinct Fourier modes,
(5.12) $A_{\mu}=\int_{H}w^{2}\,dV_{H}=\operatorname{vol}(H)+\frac{1}{2}\,\operatorname{vol}(H)\sum_{m\geq 1}a_{m}^{2}.$
Moreover, for $\rho\in(0,1/10]$,
$\frac{F(\rho)}{\operatorname{vol}(B^{s})\rho^{s}}$ $=\int_{H}w\,\mathcal{A}_{\rho}w\,dV_{H}$
(5.13) $=\operatorname{vol}(H)+\frac{1}{2}\,\operatorname{vol}(H)\sum_{m\geq 1}a_{m}^{2}\,m_{s}(2\pi|k_{m}|\rho).$
Therefore
(5.14) $q(\rho)=\frac{F(\rho)}{\operatorname{vol}(B^{s})A_{\mu}\,\rho^{s}}=\frac{1+\frac{1}{2}\sum_{m\geq 1}a_{m}^{2}\,m_{s}(2\pi|k_{m}|\rho)}{1+\frac{1}{2}\sum_{m\geq 1}a_{m}^{2}}.$
Since $\sum a_{m}^{2}<\infty$ and $|m_{s}|\leq 1$, dominated convergence and $m_{s}(0)=1$ give $q(\rho)\to 1$ as $\rho\to 0^{+}$.
Choose numbers $0<\varepsilon_{0}<1<L_{0}$ such that
(5.15) $|m_{s}(\tau)-1|\leq\tfrac{1}{100}\quad\text{for }0\leq\tau\leq\varepsilon_{0},\qquad|m_{s}(\tau)|\leq\tfrac{1}{100}\quad\text{for }\tau\geq L_{0}.$
(This is possible by continuity at 0 and $\lim_{\tau\to\infty}m_{s}(\tau)=0$.)
For each $m$ set
$\eta_{m}:=\frac{1}{100}\,a_{m}^{2},$
and let $U_{m}:=U(\eta_{m})\geq 2$ be the scale parameter given by Lemma 5.2. Define
$r_{m}:=\frac{L_{0}}{2\pi|k_{m}|}\,.$
We choose the frequencies inductively so that $|k_{m}|$ is strictly increasing and
(5.16) $2\pi|k_{m-1}|\,U_{m}r_{m}\leq\varepsilon_{0},\qquad U_{m}r_{m}\leq\tfrac{1}{10},\qquad r_{m}^{\delta}\leq\frac{1}{10m}\,\eta_{m}.$
This is possible because once $a_{m}$ (hence $\eta_{m}$ and $U_{m}$) is fixed, we are free to take $|k_{m}|$ arbitrarily large, which makes $r_{m}$ arbitrarily small.
Fix $m$ and take any $\rho\in[r_{m},U_{m}r_{m}]$. Then
$2\pi|k_{m}|\rho\geq 2\pi|k_{m}|r_{m}=L_{0},$
so (5.15) gives $|m_{s}(2\pi|k_{m}|\rho)|\leq 1/100$, and the same bound holds for all $j\geq m$ since $|k_{j}|\geq|k_{m}|$. On the other hand, (5.16) implies, for $j<m$,
$2\pi|k_{j}|\rho\leq 2\pi|k_{m-1}|\,U_{m}r_{m}\leq\varepsilon_{0},$
so (5.15) gives $m_{s}(2\pi|k_{j}|\rho)\geq 1-1/100$ for all $j<m$.
Let $A:=\sum_{j<m}a_{j}^{2}$ and $T:=\sum_{j\geq m}a_{j}^{2}$ (so $T\geq a_{m}^{2}$). Using the bounds above in (5.14) yields, for $\rho\in[r_{m},U_{m}r_{m}]$,
$q(\rho)\leq\frac{1+\frac{1}{2}\bigl(A+\tfrac{1}{100}T\bigr)}{1+\frac{1}{2}(A+T)}\leq 1-\frac{1-\tfrac{1}{100}}{1+\frac{1}{2}(A+T)}\cdot\frac{T}{2}.$
Since $\sum_{m\geq 1}a_{m}^{2}\leq\sum_{m\geq 1}a_{m}\leq\tfrac{1}{4}$, we have $A+T\leq\tfrac{1}{4}$ and hence $1+\frac{1}{2}(A+T)\leq\tfrac{9}{8}$. Therefore
$q(\rho)\leq 1-\frac{4}{9}\Bigl(1-\tfrac{1}{100}\Bigr)T\leq 1-\eta_{m},$
so $q(\rho)\leq 1-\eta_{m}$ on the entire block $r_{m},U_{m}r_{m}]$.
Applying Lemma 5.2 with $r=r_{m}$ and $\eta=\eta_{m}$ gives
$\left|\frac{G(r_{m}^{2})}{(r_{m}^{2})^{s/2}}-\gamma_{s}A_{\mu}\right|=\left|\frac{G(r_{m}^{2})}{r_{m}^{s}}-\gamma_{s}A_{\mu}\right|\geq c_{s}\,A_{\mu}\,\eta_{m}.$
Finally, the last condition in (5.16) implies
$\frac{\eta_{m}}{(r_{m}^{2})^{\delta/2}}=\frac{\eta_{m}}{r_{m}^{\delta}}\geq 10m\longrightarrow\infty,$
so no estimate of the form $\bigl|G(t)/t^{s/2}-\gamma_{s}A_{\mu}\bigr|=O(t^{\delta/2})$ can hold as $t\to 0^{+}$. By the implication (5.8), this contradicts the consequence of a uniform remainder $O(\lambda^{n-s-\delta})$. Therefore no uniform power saving $O(\lambda^{n-s-\delta})$ can hold over our class, and the $o(\lambda^{n-s})$ remainder in Theorem 1.3 is essentially sharp.

### v2

@@
 $N_{\mu}(\lambda):=\sum_{\lambda_{j}\leq\lambda}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}.$
-Assume that $\mu$ is $s$-Ahlfors regular for some $s\in(0,n)$ and admits an averaged $s$-density constant $A_{\mu}$. We prove that
+Assume that $\mu$ admits an averaged $s$-density constant $A_{\mu}$ with correlation dimension $s\in(0,n)$. We prove that
 $N_{\mu}(\lambda)=(2\pi)^{-(n-s)}\,\operatorname{vol}(B^{\,n-s})\,A_{\mu}\,\lambda^{n-s}+o(\lambda^{n-s})\qquad(\lambda\to\infty).$
-The hypotheses of $s$-Ahlfors regularity and the averaged $s$-density condition are essentially optimal for such a one-term asymptotic, and in general the remainder $o(\lambda^{n-s})$ cannot be improved uniformly to a power-saving error term.
-This extends the classical Kuznecov formulae of Zelditch [Zel92] for smooth submanifold measures to a broad class of singular and fractal measures.
+The averaged $s$-density condition is necessary for such a one-term asymptotic, and in general, the remainder $o(\lambda^{n-s})$ is sharp in the sense that it cannot be improved uniformly to a power-saving error term.
+This extends the classical Kuznecov formula of Zelditch [Zel92] for smooth submanifold measures to a broad class of singular and fractal measures.
 1. Introduction
@@
 (1.2) $\sum_{\lambda_{j}\leq\lambda}\biggl|\int_{H}e_{j}\,dV_{H}\biggr|^{2}=(2\pi)^{-(n-k)}\,\operatorname{vol}(B^{n-k})\,\operatorname{vol}(H)\,\lambda^{\,n-k}+O(\lambda^{n-k-1}),$
-where $B^{n-k}\subset\mathbb{R}^{n-k}$ is the Euclidean unit ball and $\operatorname{vol}(H)$ denotes the kk–dimensional Riemannian volume of $H$.
+where $B^{n-k}\subset\mathbb{R}^{n-k}$ is the Euclidean unit ball and $\operatorname{vol}(H)$ denotes the $k$-dimensional Riemannian volume of $H$.
 Thus the leading growth of the Kuznecov sum is governed by the ambient dimension $n$, the submanifold dimension $k$, and an explicit geometric factor.
@@
 Throughout the paper we use the following convention for Euclidean ball volumes.
-For $m>0$ (not necessarily an integer), we write $\operatorname{vol}(B^{m})$ for the “mm–dimensional volume” of the Euclidean unit ball in formal dimension $m$, defined by
-$\operatorname{vol}(B^{m}):=\frac{\pi^{m/2}}{\Gamma\bigl(\frac{m}{2}+1\bigr)}.$
-When $m\in\mathbb{N}$, this coincides with the usual Lebesgue mm–volume of the unit ball in $\mathbb{R}^{m}$.
-For an embedded kk–dimensional submanifold $H\subset M$ we write $\operatorname{vol}(H)$ for its kk–dimensional Riemannian volume.
-We now describe the small-scale geometric hypotheses on $\mu$ that will play the role of “dimension” and “density” in our setting.
+For $s\in\mathbb{R}^{+}$, we write $\operatorname{vol}(B^{s})$ for the “$s$-dimensional volume” of the Euclidean unit ball in formal dimension $s$, defined by
+$\operatorname{vol}(B^{s}):=\frac{\pi^{s/2}}{\Gamma\bigl(\frac{s}{2}+1\bigr)}.$
+When $s\in\mathbb{Z}^{+}$, this coincides with the usual Lebesgue $s$-volume of the unit ball in $\mathbb{R}^{s}$.
+For an embedded $k$-dimensional submanifold $H\subset M$ we write $\operatorname{vol}(H)$ for its $k$-dimensional Riemannian volume.
+We now describe the small-scale hypotheses on $\mu$ that will play the role of “dimension” and “volume” in our setting.
 Throughout, $B(x,r)$ denotes the geodesic ball of radius $r$ centered at $x\in M$ with respect to the Riemannian distance $d_{g}$.
-Definition 1.1 ($s$-Ahlfors regular measure).
-Let $s\in(0,n)$. A finite Borel measure $\mu$ on $M$ is called $s$-Ahlfors regular if there exist constants $c,C>0$ and $r_{0}>0$ such that
-(1.3) $c\,r^{s}\leq\mu(B(x,r))\leq C\,r^{s},\qquad x\in\operatorname{supp}\mu,\ 0<r\leq r_{0}.$
-This condition is standard in geometric measure theory and implies that the Hausdorff dimension of $\mu$ is $s$.
-See, for example, [Mat15].
-It is satisfied by Riemannian volume measures on $s$-dimensional embedded submanifolds, as well as by many self-similar and self-conformal measures on fractals.
-To obtain a one-term asymptotic for (1.1), one needs a stabilization of the averaged small-scale behavior of $\mu$. For $r\geq 0$ define the distance distribution
+For $r\geq 0$ define the distance distribution
 (1.3) $F(r):=(\mu\times\mu)\bigl(\{(x,y)\in M\times M:\ d_{g}(x,y)\leq r\}\bigr)=\int_{M}\mu\bigl(B(x,r)\bigr)\,d\mu(x).$
+The small-scale growth rate of $F(r)$ encodes the pair statistics of $\mu$ and leads naturally to the correlation dimension of $\mu$, see for instance [MMR00].
+Definition 1.1 (Correlation dimension).
+Let $\mu$ be a finite Borel measure on $M$. If the following limit exists, we define the correlation dimension of $\mu$ to be
+(1.4) $\dim_{C}(\mu):=\lim_{r\to 0^{+}}\frac{\log F(r)}{\log r}.$
+To obtain a one-term asymptotic for (1.1), it is necessary to assume that the small-scale behavior of $F(r)$ stabilizes in the following sense.
 Definition 1.2 (Averaged $s$-density).
-Let $s\in(0,n)$. We say that $\mu$ admits an averaged $s$-density if there exists a constant $A_{\mu}\in(0,\infty)$ (called the averaged $s$-density constant of $\mu$) such that
+Let $s\in(0,n)$. We say that $\mu$ admits an averaged $s$-density constant $A_{\mu}\in(0,\infty)$ if
 (1.5) $F(r)=\operatorname{vol}(B^{s})\,A_{\mu}\,r^{s}+o(r^{s})\qquad(r\to 0^{+}).$
@@
 (1.6) $A_{\mu}=\lim_{r\to 0^{+}}\frac{F(r)}{\operatorname{vol}(B^{s})\,r^{s}}=\lim_{r\to 0^{+}}\frac{1}{\operatorname{vol}(B^{s})\,r^{s}}\int_{M}\mu(B(x,r))\,d\mu(x).$
+When (1.5) holds, it forces $\dim_{C}(\mu)=s$, and $A_{\mu}$ is the corresponding averaged $s$-dimensional density constant arising from the distance distribution.
 We can now state our main result.
-It shows that under the Ahlfors regularity assumption and the averaged $s$-density condition (1.5), the Kuznecov sum (1.1) has a one-term asymptotic with exponent nsn-s and an explicit leading constant depending only on $n$ and $s$.
+It shows that once the distance distribution $F(r)$ admits a stabilized small-scale asymptotic in the sense of (1.5), the Kuznecov sum (1.1) has a one-term asymptotic with exponent nsn-s and an explicit leading constant.
 Theorem 1.3.
 Let $(M,g)$ be a compact, connected Riemannian manifold of dimension $n\geq 2$.
-Fix $s\in(0,n)$ and let $\mu$ be a finite Borel measure on $M$ which is $s$-Ahlfors regular in the sense of (1.3) and admits an averaged $s$-density constant $A_{\mu}$ in the sense of (1.5).
-Then the Kuznecov sum (1.1) satisfies
+Fix $s\in(0,n)$ and let $\mu$ be a finite Borel measure on $M$ admitting an averaged $s$-density constant $A_{\mu}$ in the sense of (1.5).
+Then
 (1.7) $N_{\mu}(\lambda)=C_{n,s}\,A_{\mu}\,\lambda^{\,n-s}+o(\lambda^{\,n-s})\qquad(\lambda\to\infty),$
@@
 The proof combines a heat-kernel regularization of the Kuznecov sum with a Tauberian argument, with the leading term determined by the small-scale distance distribution of $\mu$.
+Combining (1.7) with Weyl’s law for the eigenvalue counting function shows that the Fourier coefficients $\int_{M}e_{j}\,d\mu$ have a quantitative mean-square decay on average.
+The precise statement is recorded in the following corollary.
+Corollary 1.4 (Average Fourier decay).
+Under the hypotheses of Theorem 1.3, we have
+$\frac{1}{\#\{j:\lambda_{j}\leq\lambda\}}\sum_{\lambda_{j}\leq\lambda}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}\approx\lambda^{-s},\qquad(\lambda\to\infty).$
+If one wishes to interpret the exponent $s$ more geometrically, a natural setting is the class of Ahlfors–David regular measures.
+Recall that a finite Borel measure $\mu$ on $M$ is $s$-Ahlfors–David regular if there exist constants $c,C>0$ and $r_{0}>0$ such that
+(1.8) $c\,r^{s}\leq\mu(B(x,r))\leq C\,r^{s},\qquad x\in\operatorname{supp}\mu,\ 0<r\leq r_{0}.$
+For such measures it is standard that the Hausdorff dimension of $\mu$ equals $s$. Thus, within the Ahlfors–David regular class, the exponent in (1.7) may be written as $n-\dim_{H}(\mu)$.
 Remark 1.5.
-The averaged $s$-density condition (1.5) is automatically satisfied under the stronger pointwise density hypothesis.
-Namely, suppose that $\mu$ is $s$-Ahlfors regular and admits a pointwise $s$-density $\theta$ in the sense that
+The averaged $s$-density condition (1.5) is automatically satisfied under a stronger pointwise density hypothesis, together with a uniform upper $s$-ball condition.
+Suppose that there exist $C>0$ and $r_{0}>0$ such that $\mu(B(x,r))\leq Cr^{s}$ for all $x\in\operatorname{supp}\mu$ and $0<r\leq r_{0}$, and that the pointwise $s$-density
 $\theta(x)=\lim_{r\to 0^{+}}\frac{\mu(B(x,r))}{\operatorname{vol}(B^{s})\,r^{s}}$
-exists for $\mu$-a.e. $x$. Then by the Ahlfors upper bound $\mu(B(x,r))/(\operatorname{vol}(B^{s})r^{s})\lesssim 1$ and dominated convergence,
-$\frac{F(r)}{\operatorname{vol}(B^{s})\,r^{s}}=\int_{M}\frac{\mu(B(x,r))}{\operatorname{vol}(B^{s})\,r^{s}}\,d\mu(x)\longrightarrow\int_{M}\theta(x)\,d\mu(x),\qquad(r\to 0^{+}).$
-Hence (1.5) holds with
-$A_{\mu}=\int_{M}\theta\,d\mu.$
+exists for $\mu$-a.e. $x$. Then dominated convergence implies
+$\frac{F(r)}{\operatorname{vol}(B^{s})\,r^{s}}=\int_{M}\frac{\mu(B(x,r))}{\operatorname{vol}(B^{s})\,r^{s}}\,d\mu(x)\longrightarrow\int_{M}\theta(x)\,d\mu(x),\qquad(r\to 0^{+}),$
+so (1.5) holds with $A_{\mu}=\int_{M}\theta\,d\mu$.
 In particular, (1.5) is strictly weaker than the existence of a pointwise $s$-density.
-The leading exponent in (1.7) is nsn-s, determined by the ambient dimension $n$ and the Hausdorff dimension $s$ of $\mu$, and the leading coefficient factors as a universal constant times the averaged $s$-density constant $A_{\mu}$.
-In the special case where $\mu=dV_{H}$ is the Riemannian volume measure on an embedded submanifold $H$ of dimension $s$, $\mu$ admits a pointwise $s$-density $\theta$ and hence satisfies the averaged $s$-density condition (1.5) by Remark 1.4.
+Indeed, by a theorem of Preiss [Pre87], if a (Radon) measure has an $s$-density $\theta$ with $0<\theta(x)<\infty$ for $\mu$-a.e.
+$x$, then necessarily $s\in\mathbb{N}$ and $\mu$ is $s$-rectifiable.
+This integer-dimensional rigidity does not apply to the averaged density condition (1.5), which allows genuinely non-integer $s$.
+See Proposition 5.3.
+In the special case where $\mu=dV_{H}$ is the Riemannian volume measure on an embedded submanifold $H\subset M$ of dimension $s$, $\mu$ admits a pointwise $s$-density $\theta$ and hence satisfies the averaged $s$-density condition (1.5) by Remark 1.5.
 In fact $\theta\equiv 1$ and $A_{\mu}=\operatorname{vol}(H)$, so (1.7) reduces to
 $\sum_{\lambda_{j}\leq\lambda}\biggl|\int_{H}e_{j}\,dV_{H}\biggr|^{2}=(2\pi)^{-(n-s)}\,\operatorname{vol}(B^{\,n-s})\,\operatorname{vol}(H)\,\lambda^{\,n-s}+o(\lambda^{\,n-s}),$
-whose leading term agrees with Zelditch’s Kuznecov formulae (1.2) with $k=s$.
-Thus Theorem 1.3 extends the classical Kuznecov formulae to a large class of singular and fractal measures.
-To the best of our knowledge, Theorem 1.3 is the first Kuznecov-type asymptotic for fractal measures.
+whose leading term agrees with Zelditch’s Kuznecov formula (1.2) with $k=s$.
+More generally, Theorem 1.3 implies that the same conclusion holds when $H$ has only weak regularity.
+For instance, suppose that $H\subset M$ is a compact embedded Lipschitz $s$-dimensional submanifold with Lipschitz boundary.
+Let $\mu$ be the $s$-dimensional Hausdorff measure induced by $d_{g}$ on $H$.
+Then $\mu$ admits a pointwise $s$-density $\theta\equiv 1$ for $\mu$-a.e.
+$x\in H$.
+Hence $\mu$ satisfies (1.5) and Theorem 1.3 yields the same leading asymptotic with $A_{\mu}=\operatorname{vol}(H)$.
+This provides a Kuznecov asymptotic at Lipschitz regularity, whereas the classical results for submanifold measures assume that $H$ is smooth and boundaryless.
+In summary, Theorem 1.3 extends the classical Kuznecov formulae to a broad class of singular and fractal measures.
+To the best of our knowledge, Theorem 1.3 provides the first Kuznecov-type asymptotic for fractal measures.
 Moreover, as discussed in [WX23a], the left-hand side of Zelditch’s Kuznecov formula (1.2) can be viewed as “audible” spectral information associated with the submanifold $H$, while the right-hand side encodes geometric information about $H$, namely its dimension $k$ and its $k$-dimensional Riemannian volume $\operatorname{vol}(H)$.
-Theorem 1.3 extends this audible–geometric correspondence from smooth submanifold measures to $s$-Ahlfors regular measures admitting an averaged $s$-density constant, so that the Hausdorff dimension $s$ and the effective “volume” $A_{\mu}$ of the measure are also audible through the leading asymptotics of $N_{\mu}(\lambda)$.
+Theorem 1.3 extends this audible–geometric correspondence to measures $\mu$ whose distance distribution admits an averaged $s$-density constant $A_{\mu}$ in the sense of (1.5).
+In this case the correlation dimension $s$ and the effective “volume” $A_{\mu}$ are audible through the leading asymptotics of $N_{\mu}(\lambda)$.
+If, in addition, $\mu$ is $s$-Ahlfors–David regular, then its Hausdorff dimension is also audible.
 For $0<u<n$ let
@@
 They established energy formulae on compact Riemannian manifolds which relate $I_{u}(\mu)$ to weighted sums of the projections of $\mu$ onto Laplace eigenspaces.
-In Section 4 we show that their identities imply the following general upper bound without assuming the existence of an averaged $s$-density constant.
+In Section 4 we show that their identities imply the following general upper bound, without assuming the existence of an averaged $s$-density constant.
 Theorem 1.6.
@@
 $\sum_{\lambda_{j}\leq\lambda}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}\leq C_{1}\,\lambda^{\,n-u}\qquad(\lambda\geq 1).$
-In particular, if $\mu$ is $s$-Ahlfors regular for some $s\in(0,n)$, then $I_{u}(\mu)<\infty$ for every $u<s$, and hence for each $\varepsilon>0$ there exists a constant $C_{2}=C_{2}(M,g,s,\varepsilon,\mu)$ such that
+In particular, if $\mu$ has correlation dimension $\dim_{C}(\mu)=s$ for some $s\in(0,n)$, then $I_{u}(\mu)<\infty$ for every $u<s$, and hence for each $\varepsilon>0$ there exists a constant $C_{2}=C_{2}(M,g,s,\varepsilon,\mu)$ such that
 $\sum_{\lambda_{j}\leq\lambda}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}\leq C_{2}\,\lambda^{\,n-s+\varepsilon}\qquad(\lambda\geq 1).$
@@
 Sharpness
-We point out that the assumptions in Theorem 1.3 and Theorem 1.5 are essentially optimal.
-In Section 5 we give examples showing, on the one hand, that one cannot dispense with the $s$-Ahlfors regularity assumption if one wishes to control the exponent purely in terms of the Hausdorff dimension of $\mu$, and on the other hand, that some stabilization of the small-scale distance distribution—such as the averaged $s$-density condition (1.5)—is needed to obtain a genuine one-term asymptotic.
-Under $s$-Ahlfors regularity alone one can at best expect upper and lower bounds of order $\lambda^{n-s}$, but not the existence of a limit for $N_{\mu}(\lambda)/\lambda^{n-s}$.
-In addition to these structural restrictions, we show in Section 5 that the remainder term in Theorem 1.3 cannot, in general, be improved to a uniform power saving over the class of measures under consideration.
-More precisely, we construct explicit examples showing that there is no exponent $\delta>0$ such that for every $s$-Ahlfors regular measure with an averaged $s$-density constant $A_{\mu}$ one has
-$N_{\mu}(\lambda)=C_{n,s}A_{\mu}\,\lambda^{n-s}+O\bigl(\lambda^{n-s-\delta}\bigr),$
-with an implicit constant depending only on $(M,g,n,s,\mu)$. The $o(\lambda^{n-s})$ remainder in (1.7) is therefore essentially sharp at this level of generality.
+We record two statements, proved in Section 5, which together explain why the averaged $s$-density hypothesis (1.5) in Theorem 1.3 is essentially optimal for a one-term asymptotic, and why the remainder $o(\lambda^{n-s})$ in (1.7) cannot be improved uniformly.
+Even for Ahlfors–David regular measures, a genuine one-term asymptotic requires stabilization of the small-scale geometry of $\mu$.
+The next proposition shows that without (1.5), a one-term asymptotic need not hold.
+Proposition 1.7 (Necessity of averaged $s$-density).
+There exists an $s$-Ahlfors–David regular measure $\mu$ for which the quantity $N_{\mu}(\lambda)/\lambda^{n-s}$ fails to converge as $\lambda\to\infty$.
+Beyond these structural issues, we also show that the remainder $o(\lambda^{n-s})$ in (1.7) is essentially optimal at this level of generality: there is no uniform power-saving improvement valid even over the Ahlfors–David regular subclass.
+Proposition 1.8 (Sharpness of the remainder).
+Fix $\delta\in(0,1)$. There exists an $s$-Ahlfors–David regular measure $\mu$ admitting an averaged $s$-density constant $A_{\mu}$ such that
+$N_{\mu}(\lambda)=C_{n,s}A_{\mu}\,\lambda^{n-s}+O(\lambda^{n-s-\delta})$
+fails to hold.
 Organization
-In Section 2 we introduce a heat-kernel regularization of the Kuznecov sum, and establish its small-time asymptotics in terms of the averaged $s$-density constant $A_{\mu}$.
+In Section 2 we introduce a heat-kernel regularization of the Kuznecov sum and establish its small-time asymptotics in terms of the averaged $s$-density constant $A_{\mu}$.
 Section 3 interprets this smoothed Kuznecov sum as a Laplace–Stieltjes transform and applies a Tauberian theorem to prove Theorem 1.3 and identify the constant $C_{n,s}$.
 In Section 4 we revisit the energy identities of Hare and Roginskaya to derive general polynomial upper bounds for $N_{\mu}(\lambda)$, proving Theorem 1.6.
-Finally, Section 5 is devoted to sharpness: we give examples showing that the Ahlfors regularity and averaged $s$-density hypotheses in Theorem 1.3 are essentially optimal, and that the $o(\lambda^{n-s})$ remainder cannot, in general, be improved to a uniform power-saving error term.
-Acknowledgement
+Finally, Section 5 is devoted to sharpness and examples: we give examples showing that the averaged $s$-density hypothesis in Theorem 1.3 is essentially optimal for a genuine one-term asymptotic, and that the $o(\lambda^{n-s})$ remainder cannot, in general, be improved uniformly to a power-saving error term.
+Acknowledgements
 This project is supported by the National Key R&D Program of China under Grant No.
@@
 LR25A010001.
+The author thanks Shi-Lei Kong for some helpful discussions.
 The author thanks Xucheng Hu and Diankun Liu for carefully reading an earlier draft of this paper.
@@
 Lemma 2.1.
-There exist $r_{0}>0$, $t_{0}>0$, and smooth functions $u_{k}(x,y)$ on $\{(x,y)\in M\times M:d_{g}(x,y)<2r_{0}\}$, $k=0,1,2,\dots$, with $u_{0}(x,x)\equiv 1$, such that for every $N\in\mathbb{N}$ we have
+There exist $r_{0}>0$, $t_{0}>0$, and smooth functions $u_{k}(x,y)$ on $\{(x,y)\in M\times M:d_{g}(x,y)<2r_{0}\}$, $k=0,1,2,\dots$, with $u_{0}(x,x)\equiv 1$, such that for every $N\in\mathbb{Z}^{+}$ we have
 (2.1) $p_{t}(x,y)=(4\pi t)^{-n/2}e^{-d_{g}(x,y)^{2}/(4t)}\biggl(\sum_{k=0}^{N-1}t^{k}u_{k}(x,y)\biggr)+R_{N}(t,x,y),$
@@
 Lemma 2.3.
-Assume that $\mu$ satisfies (1.3) and (1.5). Then, as $t\to 0^{+}$,
+Assume that $\mu$ satisfies (1.5). Then, as $t\to 0^{+}$,
 (2.6) $G(t)=\gamma_{s}\,A_{\mu}\,t^{s/2}+o(t^{s/2}),$
@@
 $F(r)=\operatorname{vol}(B^{s})A_{\mu}r^{s}+\mathcal{E}(r),$
-with $\mathcal{E}(r)=o(r^{s})$ as $r\to 0^{+}$. Fix $r_{0}>0$ as in (1.3). Split
+with $\mathcal{E}(r)=o(r^{s})$ as $r\to 0^{+}$. Fix $r_{0}>0$ to be a small constant. Split
 $G(t)=\int_{0}^{r_{0}}e^{-r^{2}/(4t)}\,dF(r)+\int_{r_{0}}^{\operatorname{diam}(M)}e^{-r^{2}/(4t)}\,dF(r)=:G_{0}(t)+G_{1}(t).$
@@
 for some $C_{1}$ depending only on $s$.
-On $\delta,r_{0}]$ we use the upper Ahlfors bound $F(r)\leq Cr^{s}$ to deduce that $|\mathcal{E}(r)|\leq C^{\prime}r^{s}$ on this interval. Thus
+On $\delta,r_{0}]$ we use the upper bound $F(r)\leq Cr^{s}$ to deduce that $|\mathcal{E}(r)|\leq C^{\prime}r^{s}$ on this interval. Thus
 $|I_{2}(t)|\leq\frac{C^{\prime}}{2t}\int_{\delta}^{r_{0}}r^{s+1}e^{-r^{2}/(4t)}\,dr\leq C^{\prime\prime}t^{-1}e^{-\delta^{2}/(4t)},$
@@
 Lemma 2.4.
-Assume that $\mu$ satisfies (1.3) and (1.5). Then
+Assume that $\mu$ satisfies (1.5). Then
 (2.9) $H_{\mu}(t)=\gamma_{n,s}\,A_{\mu}\,t^{-(n-s)/2}+o\bigl(t^{-(n-s)/2}\bigr)\qquad(t\to 0^{+}),$
@@
 as $t\to 0^{+}$.
-For the second integral we use (2.3) and the Ahlfors bound via $F(r)$:
+For the second integral we use (2.3):
 $\begin{split}\biggl|\iint_{d_{g}(x,y)\leq r_{0}}e^{-d_{g}(x,y)^{2}/(4t)}\omega(x,y)\,d\mu(x)\,d\mu(y)\biggr|&\leq C_{2}\iint_{d_{g}(x,y)\leq r_{0}}d_{g}(x,y)e^{-d_{g}(x,y)^{2}/(4t)}\,d\mu(x)\,d\mu(y)\\ &=C_{2}\int_{0}^{r_{0}}re^{-r^{2}/(4t)}\,dF(r)\\ &=:C_{2}J(t).\end{split}$
@@
 so
-$(2\pi)^{-(n-s)}\,\operatorname{vol}(B^{n-s})=2^{-(n-s)}\pi^{-(n-s)}\cdot\frac{\pi^{(n-s)/2}}{\Gamma\bigl(\frac{n-s}{2}+1\bigr)}=2^{-(n-s)}\pi^{-(n-s)/2}\frac{1}{\Gamma\bigl(\frac{n-s}{2}+1\bigr)}=C_{n,s}.$
+$(2\pi)^{-(n-s)}\,\operatorname{vol}(B^{n-s})=2^{-(n-s)}\pi^{-(n-s)}\cdot\frac{\pi^{(n-s)/2}}{\Gamma\bigl(\frac{n-s}{2}+1\bigr)}=C_{n,s}.$
 Thus
@@
 (4.1) $c_{n,u}^{-1}I_{u}(\mu)\leq\sum_{j=0}^{\infty}(1+\lambda_{j}^{2})^{-(n-u)/2}\Bigl|\int_{M}e_{j}\,d\mu\Bigr|^{2}\leq c_{n,u}\,I_{u}(\mu).$
-Remark 4.2.
-Estimates of the form (4.1) are proved in [HR03] by expressing suitable powers of the Laplace–Beltrami operator as pseudodifferential operators and identifying their kernels with Riesz-type potentials.
-In particular, $I_{u}(\mu)$ is comparable, up to multiplicative constants depending only on $(M,g,n,u)$, to the quadratic form associated with the operator $(I-\Delta_{g})^{-(n-u)/2}$ evaluated on $\mu$.
-See [HR03] for details.
 Given Proposition 4.1, the proof of Theorem 1.6 is straightforward.
@@
 This proves the first assertion of Theorem 1.6.
-To obtain the “in particular” statement, it remains to note that if $\mu$ is $s$-Ahlfors regular for some $s\in(0,n)$, then $I_{u}(\mu)<\infty$ for every $u<s$.
-This is a standard estimate, but we include a brief argument in our setting for completeness.
+To obtain the second statement of Theorem 1.6, it remains to note that if $\mu$ has correlation dimension $\dim_{C}(\mu)=s$ for some $s\in(0,n)$, then $I_{u}(\mu)<\infty$ for every $u<s$.
+This is a standard estimate, and is the reason why the correlation dimension is sometimes called the energy dimension.
+We include a brief argument in our setting for completeness.
 Lemma 4.2.
-Let $\mu$ be an $s$-Ahlfors regular measure on $M$ with $s\in(0,n)$, and let $u\in(0,s)$. Then $I_{u}(\mu)<\infty$.
+Let $\mu$ be a finite Borel measure on $M$ with correlation dimension $\dim_{C}(\mu)=s$ for some $s\in(0,n)$, and let $u\in(0,s)$. Then $I_{u}(\mu)<\infty$.
 Proof.
-Recall
+Recall that
 $F(r)=(\mu\times\mu)\bigl(\{(x,y):d_{g}(x,y)\leq r\}\bigr)=\int_{M}\mu(B(x,r))\,d\mu(x).$
 Using the layer-cake representation, we have
-$I_{u}(\mu)=\int_{0}^{\operatorname{diam}(M)}ur^{-u-1}F(r)\,dr+\mu(M)^{2}\,\operatorname{diam}(M)^{-u},$
+$I_{u}(\mu)=\int_{0}^{\operatorname{diam}(M)}ur^{-u-1}F(r)\,dr+\mu(M)^{2}\,\operatorname{diam}(M)^{-u}.$
+Fix $\varepsilon>0$ so that $u<s-\varepsilon$. Since $\dim_{C}(\mu)=s$, there exist constants $C>0$ and $r_{0}>0$ such that
+$F(r)\leq Cr^{s-\varepsilon}\qquad(0<r\leq r_{0}).$
 Split the integral at $r_{0}$:
@@
 On $(0,r_{0}]$,
-$u\int_{0}^{r_{0}}r^{-u-1}F(r)\,dr\leq uC\int_{0}^{r_{0}}r^{s-u-1}\,dr=\frac{uC}{s-u}\,r_{0}^{\,s-u}<\infty,$
-since su>0s-u>0. On $r_{0},\operatorname{diam}(M)]$,
+$u\int_{0}^{r_{0}}r^{-u-1}F(r)\,dr\leq uC\int_{0}^{r_{0}}r^{s-\varepsilon-u-1}\,dr<\infty,$
+since $s-\varepsilon-u>0$. On $r_{0},\operatorname{diam}(M)]$,
 $u\int_{r_{0}}^{\operatorname{diam}(M)}r^{-u-1}F(r)\,dr\leq u\mu(M)^{2}\int_{r_{0}}^{\operatorname{diam}(M)}r^{-u-1}\,dr<\infty.$
@@
 Remark 4.3.
-Theorem 1.5 shows that for each fixed $u<s$ one can control the growth of the Kuznecov sum by $\lambda^{n-u}$ under the mild hypothesis $I_{u}(\mu)<\infty$, and in particular for $s$-Ahlfors regular measures.
+Theorem 1.6 shows that for each fixed $u<s$ one can control the growth of the Kuznecov sum by $\lambda^{n-u}$ under the mild hypothesis $I_{u}(\mu)<\infty$.
 Our main theorem, Theorem 1.3, goes further by identifying the sharp exponent nsn-s and the exact leading constant in the asymptotic (1.7), under the additional assumption of the averaged $s$-density condition.
@@
 This is why we turn to a more refined analysis based on the heat kernel and Tauberian theory.
-5. Sharpness of Theorem 1.3
-In this section, we discuss the sharpness of Theorem 1.3.
-We explain the necessity of the two structural hypotheses, namely $s$-Ahlfors regularity and the averaged $s$-density condition (1.5), for obtaining the exponent nsn-s and a genuine one-term asymptotic.
-We also show that the remainder $o(\lambda^{n-s})$ in (1.7) is essentially optimal at this level of generality: in general it cannot be improved uniformly to a power-saving error term over the class of measures under consideration.
-5.1. Necessity of the Ahlfors regularity assumption
-It is natural to ask whether the exponent nsn-s in Theorem 1.3 (or in the upper bound of Theorem 1.5) could be expressed solely in terms of the Hausdorff dimension of $\mu$, without any Ahlfors regularity hypothesis.
-This is not possible in general.
-Let $H_{1},H_{2}\subset M$ be disjoint embedded submanifolds with dimensions $s_{1}<s_{2}$, and let $\mu_{i}=dV_{H_{i}}$ and $\mu=\mu_{1}+\mu_{2}$. Then
-$\dim_{H}(\mu)=\max(s_{1},s_{2})=s_{2},$
-but the growth of $N_{\mu}(\lambda)$ is governed by the smaller dimension $s_{1}$.
-To see this, write
-$a_{j}:=\int_{H_{1}}e_{j}\,dV_{H_{1}},\qquad b_{j}:=\int_{H_{2}}e_{j}\,dV_{H_{2}}.$
-Then
-$N_{\mu}(\lambda)=\sum_{\lambda_{j}\leq\lambda}|a_{j}+b_{j}|^{2}=\sum_{\lambda_{j}\leq\lambda}|a_{j}|^{2}+\sum_{\lambda_{j}\leq\lambda}|b_{j}|^{2}+2\Re\sum_{\lambda_{j}\leq\lambda}a_{j}\overline{b_{j}}.$
-By Cauchy–Schwarz,
-$\Bigl|\sum_{\lambda_{j}\leq\lambda}a_{j}\overline{b_{j}}\Bigr|\leq\Bigl(\sum_{\lambda_{j}\leq\lambda}|a_{j}|^{2}\Bigr)^{1/2}\Bigl(\sum_{\lambda_{j}\leq\lambda}|b_{j}|^{2}\Bigr)^{1/2}.$
-Hence we have the robust comparison
-(5.1) $\bigl|N_{\mu}(\lambda)-N_{\mu_{1}}(\lambda)-N_{\mu_{2}}(\lambda)\bigr|\leq 2\,\bigl(N_{\mu_{1}}(\lambda)N_{\mu_{2}}(\lambda)\bigr)^{1/2}.$
-By the smooth submanifold Kuznecov formulae,
-$N_{\mu_{i}}(\lambda)\sim C_{n,s_{i}}\,\operatorname{vol}(H_{i})\,\lambda^{n-s_{i}}\qquad(i=1,2),$
-so $N_{\mu_{2}}(\lambda)=o(N_{\mu_{1}}(\lambda))$ and the right-hand side of (5.1) is
-$O\Bigl(\lambda^{\frac{n-s_{1}}{2}}\,\lambda^{\frac{n-s_{2}}{2}}\Bigr)=O\bigl(\lambda^{\,n-\frac{s_{1}+s_{2}}{2}}\bigr)=o\bigl(\lambda^{n-s_{1}}\bigr).$
-Therefore $N_{\mu}(\lambda)=N_{\mu_{1}}(\lambda)+o(\lambda^{n-s_{1}})$, and in particular
-$N_{\mu}(\lambda)\sim C_{n,s_{1}}\,\operatorname{vol}(H_{1})\,\lambda^{n-s_{1}}\qquad(\lambda\to\infty),$
-even though $\dim_{H}(\mu)=s_{2}$. Thus, no exponent depending only on $\dim_{H}(\mu)$ can govern $N_{\mu}(\lambda)$ for arbitrary finite Borel measures.
+5. Sharpness and examples
+In this section, we discuss the sharpness of Theorem 1.3 and provide a few examples.
+We explain the necessity of the averaged $s$-density condition (1.5) for obtaining a genuine one-term asymptotic, and we show that the remainder $o(\lambda^{n-s})$ in (1.7) is essentially optimal at this level of generality.
+In particular, it cannot be improved uniformly to a power-saving error term, even over natural regular subclasses of measures.
+In the final part of this section we present an explicit self-similar fractal measure whose distance distribution $F(r)$ admits an averaged $s$-density constant, and hence satisfies (1.5).
 5.1. Necessity of the averaged $s$-density condition
-The averaged $s$-density hypothesis in Theorem 1.3 is not merely a technical convenience: without a stabilization of the small-scale behavior of $\mu$, one cannot in general expect the normalized quantities $N_{\mu}(\lambda)/\lambda^{n-s}$ to converge.
+In this subsection we prove Proposition 1.7, which shows that the averaged $s$-density hypothesis in Theorem 1.3 is not merely a technical convenience.
+Without a stabilization of the small-scale behavior of $\mu$, one cannot in general expect the normalized quantities $N_{\mu}(\lambda)/\lambda^{n-s}$ to converge.
 Recall the distance distribution
@@
 $a(r):=\frac{F(r)}{r^{s}}\qquad(r>0).$
-If $\mu$ is $s$-Ahlfors regular then $a(r)$ is bounded above and below by positive constants for $0<r\leq r_{0}$.
-Since $F(0)=0$, integration by parts gives, modulo a boundary term that is exponentially decaying in $1/t$, we have
+If $\mu$ is $s$-Ahlfors–David regular then $a(r)$ is bounded above and below by positive constants for $0<r\leq r_{0}$.
+Since $F(0)=0$, an integration by parts yields, up to a boundary term that decays exponentially in $1/t$,
 (5.1) $G(t)=\frac{1}{2t}\int_{0}^{\operatorname{diam}(M)}re^{-r^{2}/(4t)}\,F(r)\,dr.$
@@
 In that setting the endpoints in (5.3) are realized along suitable subsequences.
-Lemma 5.1 (Realizing $\liminf/\limsup$ from symmetric scale blocks).
+Lemma 5.1 (Realizing $\liminf/\limsup$).
 Assume $a$ is bounded on $(0,\operatorname{diam}(M)]$, and let
@@
 5.2. Optimality of the remainder
-In the smooth submanifold case, Zelditch’s Kuznecov formulae (1.2) come with an $O(\lambda^{n-k-1})$ remainder term.
+In the smooth submanifold case, Zelditch’s Kuznecov formula (1.2) comes with an $O(\lambda^{n-k-1})$ remainder term.
 Moreover, for generic metrics this remainder can be improved to $o(\lambda^{n-k-1})$ by a recent result of Kaloshin, Wyman and the author [KWX25].
 In contrast, the asymptotic formula in Theorem 1.3 is stated with an $o(\lambda^{n-s})$ remainder.
-It is natural to ask whether, under the assumptions of $s$-Ahlfors regularity and the existence of an averaged $s$-density constant, one can improve this to a uniform power-saving error term of the form
+It is natural to ask whether one can strengthen (1.7) to a uniform power-saving error term of the form
 (5.4) $N_{\mu}(\lambda)=C_{n,s}A_{\mu}\,\lambda^{n-s}+O(\lambda^{n-s-\delta})\qquad(\lambda\to\infty),$
-for some $\delta>0$ depending only on $(M,g,n,s)$ and valid for all measures $\mu$ in our class. We now explain why such a uniform improvement is not available in general.
+for some $\delta>0$ depending only on $(M,g,n,s)$ and valid for all measures $\mu$ satisfying the hypotheses of Theorem 1.3.
+Proposition 1.8 asserts that such a uniform improvement is not available in general, even under additional regularity assumptions, and we prove it in this subsection.
 Step 1: from $N_{\mu}$ to the Laplace–Stieltjes transform $H_{\mu}$. Write $\mathcal{N}(S):=N_{\mu}(\sqrt{S})$ as in (3.1), so that
@@
 $w(x):=1+\sum_{m\geq 1}a_{m}\,\phi_{m}(x),\qquad\text{and}\qquad d\mu:=w\,dV_{H}.$
-Then $\tfrac{3}{4}\leq w\leq\tfrac{5}{4}$, so $\mu$ is $s$-Ahlfors regular.
+Then $\tfrac{3}{4}\leq w\leq\tfrac{5}{4}$, so $\mu$ is $s$-Ahlfors–David regular.
 Moreover, since $\mu$ is absolutely continuous with respect to $dV_{H}$ with bounded density $w$, $\mu$ admits a pointwise $s$-density $\theta=w$.
@@
 Therefore no uniform power saving $O(\lambda^{n-s-\delta})$ can hold over our class, and the $o(\lambda^{n-s})$ remainder in Theorem 1.3 is essentially sharp.
+5.3. A self-similar example on the unit interval
+We record an explicit self-similar fractal measure on $[0,1]$ of non-integer dimension satisfying the averaged $s$-density condition (1.5).
+Proposition 5.3.
+Let $\phi_{0},\phi_{1}:[0,1]\to[0,1]$ be
+$\phi_{0}(x)=\tfrac{1}{2}x,\qquad\phi_{1}(x)=\tfrac{2}{3}+\tfrac{1}{3}x,$
+and let $K\subset[0,1]$ be the attractor $K=\phi_{0}(K)\cup\phi_{1}(K)$.
+Let $s\in(0,1)$ be the unique solution of $2^{-s}+3^{-s}=1$, and set $p_{0}:=2^{-s}$, $p_{1}:=3^{-s}$.
+Let $\mu$ be the unique Borel probability measure supported on $K$ satisfying
+(5.16) $\mu(A)=p_{0}\,\mu(\phi_{0}^{-1}(A))+p_{1}\,\mu(\phi_{1}^{-1}(A))\qquad\text{for all Borel }A\subset[0,1].$
+Then $s\notin\mathbb{N}$, and $\mu$ admits an averaged $s$-density constant $A_{\mu}\in(0,\infty)$ in the sense of (1.5).
+Proof.
+It is clear that $s$ is not an integer. For $0<r<1/6$, pairs from different first-level pieces cannot satisfy $|x-y|\leq r$, hence expanding $\mu\times\mu$ using (5.16) gives the exact recursion
+$F(r)=p_{0}^{2}\,F(2r)+p_{1}^{2}\,F(3r)\qquad(0<r<1/6).$
+Setting $f(r):=F(r)/(\operatorname{vol}(B^{s})\,r^{s})$ and using $p_{0}=2^{-s}$, $p_{1}=3^{-s}$ yields
+$f(r)=p_{0}\,f(2r)+p_{1}\,f(3r)\qquad(0<r<1/6).$
+With $t=-\log r$ and $g(t):=f(e^{-t})$, we obtain
+$g(t)=p_{0}\,g(t-\log 2)+p_{1}\,g(t-\log 3)\qquad(t>\log 6).$
+Since $\log 2/\log 3\notin\mathbb{Q}$, the key renewal theorem (see, e.g.
+[Fel71]) implies that $g(t)\to A_{\mu}$ as $t\to\infty$ for some $A_{\mu}\in(0,\infty)$, i.e.
+$f(r)\to A_{\mu}$ as $r\to 0^{+}$, which is exactly (1.5) with exponent $s$.
+Finally, we remark that this measure $\mu$ in our example satisfies Ahlfors–David regularity.
+∎